### v1

# Stochastic dynamic programming under recursive Epstein-Zin preferences

Abstract This paper treats of discrete-time Markov decision processes with recursive utility defined using a non-linear aggregator of Epstein-Zin type with the Kreps-Porteus certainty equivalent operator. According to the classification introduced by Marinacci and Montrucchio, the aggregators that we consider are of Thompson type. We focus on the existence and uniqueness of a solution to the Bellman equation. If the action spaces of an agent are singletons, the problem concerns the existence and uniqueness of a recursive utility in the sense of Koopmans. The models of this type were recently studied by Ren and Stachurski who applied Du’s fixed point theorem for increasing and convex or concave operators acting on an ordered Banach space. We show that, in two cases considered by Ren and Stachurski, the existence and uniqueness of solution to the Bellman equation follows from the Banach contraction mapping principle. We allow the per-period utilities in the models to be unbounded. Therefore, we work with the weighted supremum norm. Since we apply the Banach fixed point theorem for contraction mappings acting on a standard complete metric space, we need not assume any boundary conditions, which are present when the Thompson metric or Du’s theorem are used. Moreover, our results give better bounds for the geometric convergence of the value iteration algorithm leading to the solution of the Bellman equation than those obtained by Du’s fixed point theorem. Some illustrative examples are given. This paper also contains also new results on the existence and uniqueness of solutions to the Bellman equation in two cases of parameters in the Epsten-Zin aggregator, where the application of Du’s theorem is not possible. Our results are based on a relatively simple reformulation of the problem exploiting the structure of Epstein-Zin preferences.
JEL classification: C61, C62, D81, E20
Keywords: Recursive utility; Thompson aggregator; Bellman equation; Dynamic programming
1 Introduction
Markov decision processes ($MDPs$) play an important role in operations research, economics, finance, engineering and computer science, see Acemoglu (2009); Bäuerle and Rieder (2011); Hernández-Lerma and Lasserre (1996, 1999); Kochenderfer (2015); Miao (2014); Sargent and Stachurski (2024); Stachurski (2009). In the classical formulation the additive utility is constructed for a $MDP$ by a simple linear aggregator of the current utility (reward) and the discounted expected continuation value. The expectation is taken with respect to some transition probability from the current state to the following one. This approach does not take into account the risk associated with state transitions. The risk aversion of an agent (decision maker) can be introduced by replacing the standard expectation by a quasi arithmetic certainty equivalent operator induced by an increasing concave utility function on the set of non-negative real numbers. One possibility is associated with using the entropic risk measure and defining the discounted utility in a multiperiod decision process in a recursive way. Here, we must mention a classical work of Hansen and Sargent (1995). Further contributions can be found in Asienkiewicz and Jaśkiewicz (2017); Bäuerle and Jaśkiewicz (2018) and a recent book of Sargent and Stachurski (2024). Related results that use other certainty equivalents and biconvergence of recursive utilities in finite horizon models can be found in Streufert (1990); Ozaki and Streufert (1996) and Bich et al. (2018). Some examples, using a a certainty equivalent induced by a power function is discussed in Miao (2014). The classical discounted utility has an important drawback. It is unable to disentangle risk attitudes from the intertemporal elasticity of substitution, which is difficult to accept in many applications to economic (finance), to consumption/investment or resource extraction models. In order to overcome this difficulty, starting from Kreps and Porteus (1978) and Epstein and Zin (1989), the more general preferences were studied by Backus et al. (2005); Borovićka and Stachurski (2020); Epstein and Zin (1989); Marinacci and Montrucchio (2010); Miao (2014); Christensen (2022); Skiadas (2009); Sargent and Stachurski (2024). In these extensions, aggregation of utilities (rewards) over time is not additive, and the standard expected value is replaced by a certainty equivalent operator.
A large part of the literature on recursive utilities has no optimisation ingredient. An attention is paid to a construction of the recursive utility defined on a sequence of utilities (rewards) (or consumption streams) using an aggregator and given certainty equivalent. The existence of the recursive utility is a fixed point problem in a suitably chosen function space. In some cases the Banach fixed point theorem is possible to apply. A first result of this type is due Stokey and Lucas (1984). However, there are many problems, where fixed point theorems for operators on partially ordered Banach spaces are applied. The reader may consult Bloise and Vailakis (2018); Bloise et al. (2024); Marinacci and Montrucchio (2010); Sargent and Stachurski (2024) for a good review of the literature on these topics. The Epstein-Zin definition of recursive utility is written in the context of a consumption model. They suggest the following preferences that are defined recursively by the $CES$ (Constant Elasticity of Substitution) aggregator:
$U_{n}:=\left[(1-\beta)C_{n}^{1-\rho}+\beta({\cal M}_{n}(U_{n+1}))^{1-\rho}\right]^{\frac{1}{1-\rho}},$
where $0<\rho\not=1,$ $\beta\in(0,1)$ is the discount factor, $(C_{n})$ is a consumption path, $U_{n}$ is the utility value of the path onward from time period $n$ and ${\cal M}_{n}$ is a certainty equivalent. It can be defined, for instance, by the power function:
${\cal M}_{n}(U_{n+1})=E_{n}(U_{n+1}^{1-\gamma})^{\frac{1}{1-\gamma}},\quad 0<\gamma\not=1.$
Here, $E_{n}$ is the conditional expectation operator with respect to the period $n$ information. This certainty equivalent is also known under the name of Kreps-Porteous operator, see Miao (2014). The value $1/\rho$ represents the elasticity of intertemporal substitution ($EIS$) between the composite good and the certainty equivalent, whereas $\gamma$ governs the level of relative risk aversion ($RRA$) with respect to atemporal gambles. As we can see from this definition, the parameters $\gamma$ and $1/\rho$ are independent, which is important in many applications. The most empirically relevant case is $\rho<\gamma,$ implying that the agent prefers early resolution of uncertainty, see Bansal and Yaron (2004); Schorfheide et al. (2018). The recursive preferences of Epstein and Zin (1989) now form a core of the quantitative asset pricing literature, also find applications ranging from optimal taxation to fiscal policy and business cycles, see Bansal and Yaron (2004), Kaplan Violante (2014) and Schorfheide et al. (2018).
The structure of non-linear recursive preferences turned out to be harder to study. It was realised that the operators involved in the construction of recursive utility are not contractive in the supremum norm. Marinacci and Montrucchio (2010) divided the Epstein-Zin aggregators into two classes. In one of them the aggregators are called of Blackwell type and the existence and uniqueness of a recursive utility can be proved by the Banach contraction mapping principle. It is connected with the work of Blackwell (1965) on discounted dynamic programming. It seems that the work of Denardo (1967) was one of the first papers that are related to dynamic programming with recursive preferences. In the second class of aggregators, called of Thompson type, the standard methods and metrics cannot be directly applied. The concave Epstein-Zin aggregators are Thompson. Marinacci and Montrucchio (2010) exploited recursive utilities derived from the second class and obtained significant results by applying the Thompson fixed point theorem for non-linear operators on ordered normed spaces with a complete and normal positive cone (consult with Thompson (1963)). The Thompson metric used in Marinacci and Montrucchio (2010) seems to be difficult to apply to dynamic programming problems considered in this paper. Theory of monotone concave and $\alpha$-concave operators was recently adapted by Bloise and Vailakis (2018) and Balbus (2020), respectively. They demonstrate new results on the existence and uniqueness of a recursive utility as well as some applications to Bellman equations in dynamic optimisation. However, their results are applicable only to certain class of Epstein-Zin aggregators.
Dynamic programming models with convex or concave aggregators were recently studied by Ren and Stachurski (2020, 2021). They obtained valuable results on the existence and uniqueness of solutions of the Bellman equation using a fixed point theorem of Du (1990) for monotone convex or concave operators. They examined Epstein-Zin preferences with Kreps-Porteus certainty equivalents and with unbounded rewards. Here, we wish to emphasise that using the Thompson metric or the fixed point theorem of Du (1990) requires some boundary conditions implying that the fixed point of the dynamic programming operator used in studying an $MDP$ in recursive formulation belongs to the interior of its domain. These boundary conditions imply some restrictions on per-period utilities (in some cases they must be above some positive number) and have an influence on the speed of convergence of iterations of dynamic programming operators to a solution of the Bellman equation.
In this paper, we study $MDPs$ involving the Epstein-Zin preferences with the Kreps-Porteus certainty equivalent. We allow the per-period utilities to be unbounded and use the weighted supremum norms. The results of Ren and Stachurski (2020, 2021) are applied to $MDPs$ with the Epstein-Zin preferences in three cases: $0<\gamma<\rho<1,$ $1<\gamma<\rho$ and $0<\gamma<1<\rho.$ We prove the existence and uniqueness of a solution to the Bellman equation in the first two cases using the Banach contraction mapping principle for mappings on the space of upper semicontinuous functions endowed with the weighted supremum norm. This method works for some class with convex aggregators that are according to Marinacci and Montrucchio (2010) of Thompson type. Using the Banach fixed point theorem and the supremum norm we get a simple estimation of the speed of geometric convergence of the value iteration algorithm to a solution to the Bellman equation. As a starting point we can choose a constant function, in one case a null function. We give two examples with bounded rewards showing that the our value iterations converge faster than in Ren and Stachurski (2020, 2021), who applied Du’s theorem. Moreover, our regularity assumptions on the per-period utility function and inequalities involving the weight function are weaker than those used in Ren and Stachurski (2020, 2021). We also show the existence and uniqueness of the solution to the Bellman equation for $MDPs$ with some specifications of parameters, in the Epstein-Zin aggregator, where the dynamic programming operators are neither convex nor concave. The fixed point theorem of Du (1990) in these cases is useless again. An application of the Banach contraction principle to dynamic programming in $MDPs$ with Epstein-Zin preferences is possible due to some simple transformations of the Bellman equations.
2 Dynamic programming model and main results
2.1 Preliminaries
[…]
$\|v\|_{\omega}:=\sup_{s\in S}\frac{|v(s)|}{\omega(s)}.$
By ${\cal C}^{\omega}$ (${\cal U}^{\omega}$) we denote the set of all continuous (upper semicontinuous) functions on $S$ for which $\|\cdot\|_{\omega}$-norm is finite and by ${\cal C}_{+}^{\omega}$ (${\cal U}_{+}^{\omega}$) we denote the set of non-negative functions in ${\cal C}^{\omega}$ (${\cal U}^{\omega}$).
Assume that $(X,d)$ is a complete metric space and $\delta\in(0,1).$ A mapping $F:X\to X$ is called a $\delta$-contraction, if $d(Fx_{1},Fx_{2})\leq\delta d(x_{1},x_{2})$ for all $x_{1},x_{2}\in X.$ By the Banach contraction mapping principle $F$ has a unique fixed point $x_{*}\in X.$ Moreover, for any element $x_{0}\in X$
$d(F^{n}x_{0},x_{*})\leq\frac{\delta^{n}}{1-\delta}d(Fx_{0},x_{0}),$ (1)
where $F^{n}$ is the $n$-th iteration of $F$, see Section 3.12 in Aliprantis and Border (2006).
Let $Z$ a separable metric space. By $\varphi$ we denote a correspondence from $S$ to $Z.$ For any $D\subset Z$ we put
$\varphi^{-1}(D):=\{s\in S:\varphi(s)\cap D\not=\emptyset\}.$
If $\varphi^{-1}(D)$ is a closed (an open) subset in $S$ for each closed (open) subset $D$ of $Z,$ then $\varphi$ is said to be upper (lower) semicontinuous. If $\varphi$ is both upper and lower semicotinuous, then it is called continuous.
2.2 A model with stationary recursive utilities
In this paper we study a dynamic decision model defined by the following objects. $(i)$ $S$ is a set of states and is assumed to be a Borel space.
$(ii)$ $A$ is the action space and is also assumed to be a Borel space.
$(iii)$ $\mathbb{K}$ is a non-empty Borel subset of $S\!\times\!A$. We assume that for each $s\in S$, the non-empty $s$-section
$A(s)=\{a\in A:(s,a)\in\mathbb{K}\}$
of $\mathbb{K}$ represents the set of actions available in state $s$.
$(iv)$ $q$ is a Feller transition probability from $\mathbb{K}$ to $S,$ i.e., the function
$(s,a)\to\int_{S}v(s^{\prime})q(ds^{\prime}|s,a)$
is continuous for any $v\in{\cal C}.$
$(v)$ $u:\mathbb{K}\to\mathbb{R}$ is a upper semicontinuous non-negative per-period utility function.
$(vi)$ $\beta\in[0,1)$ is a discount factor.
We first study a stationary model in which an agent uses the same decision rule in every period. Let $\Phi$ be the set of all Borel measurable mappings $f$ from $S$ into $A$ such that $f(s)\in A(s)$ for each $s\in S$. By Theorem 1 in Brown and Purves (1973), $\Phi\not=\emptyset.$ A stationary policy is a constant sequence $\pi=(f,f,\ldots)$ where $f\in\Phi$ and therefore can be identified with the Borel mapping $f.$
Assume for a moment that $u$ is bounded. A state-action aggregator $H$ maps feasible state-action pairs $(s,a)\in\mathbb{K}$ and bounded Borel measurable functions $w$ into real values $H(s,a,w)$ representing lifetime utility, contingent on current action $a$, current state $s$ and the use of $w$ to evaluate future states. Traditional additively separable case is implemented by setting:
$H(s,a,w)=u(s,a)+\beta\int_{S}w(s^{\prime})q(ds^{\prime}|s,a),$
where $(s,a)\in\mathbb{K}.$ Under Epstein-Zin preferences, the state-action aggregator $H$ is of the form
$H(s,a,w)=\left[r(s,a)+\beta\left(\int_{S}w(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},\ \ (s,a)\in\mathbb{K},$ (2)
[…]
$r(s,a)=(1-\beta)u(s,a)^{1-\rho},\quad(s,a)\in\mathbb{K}.$
Definition 1
A function $\widehat{v}$ is a solution to the Bellman equation, if
$\widetilde{v}(s)=\max_{a\in A(s)}H(s,a,\widetilde{v})\quad\mbox{for all}\quad s\in S.$ (3)
Definition 2
For any policy $f\in\Phi$, we call $v_{f}$ an $f$-value function, if it satisfies the equation
$v_{f}(s)=H(s,f(s),v_{f}),\quad s\in S$
with $H$ defined in (2).
The value function $v_{f}$ is a recursive utility in the infinite horizon model in the sense of Koopmans (1960), see also Marinacci and Montrucchio (2010). It is determined by the aggregator (2) and a policy $f\in\Phi.$
Definition 3
A stationary policy $f_{*}\in\Phi$ is optimal, if
$v_{f_{*}}(s)=\widetilde{v}(s)\geq v_{f}(s)\quad\mbox{for all }f\in\Phi\mbox{ and }s\in S,$
and $\widetilde{v}$ is a solution to the Bellman equation (3).
Such a restriction to the class of stationary policies and the recursive utilities in the sense of Koopmans (1960) is also used by Ren and Stachurski (2020, 2021).
If there exists a policy $f\in\Phi$ that maximises the right-hand side of (3), then, as we show in the sequel, $f$ turns out to be an optimal stationary policy. This case is well-known in the literature with standard affine aggregators and expected value as certainty equivalent, see Bertsekas (2022); Hernández-Lerma and Lasserre (1996).
2.3 Main results on optimal policies in stationary models
The Bellman equation in (3) is precisely of the form
$\widetilde{v}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}\widetilde{v}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},$ (4)
for all $s\in S,$ provided that the integral of $\widetilde{v}(s)^{1-\gamma}$ exists with respect to $q(\cdot|s,a)$ for every $(s,a)\in\mathbb{K}.$ Define
$\theta:=\frac{1-\gamma}{1-\rho}.$
This number will be used in all four cases of parametrisation in the Epstein-Zin preferences considered in this paper. Let $\hat{0}$ be the null function.
In order to find a solution $v$ to the Bellman equation (4) when $\gamma,\ \rho\in(0,1)$, we accept the following conditions.
  • (A1)
    For each $s\in S$, the set $A(s)$ is compact and, moreover, the correspondence $s\to A(s)$ is upper semicontinuous.
  • (A2)
    There exists a continuous function $\omega:S\to[1,\infty)$ such that
    $r(s,a)\leq M\omega(s),\quad(s,a)\in\mathbb{K}$
    for some constant $M>0.$
  • (A3)
    The function
    $(s,a)\to\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)$
    is continuous and there is a constant $c\geq 1$ such that
    $\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)\leq c\omega(s)$
    for all $(s,a)\in{\mathbb{K}}$ and $c\beta^{\theta}<1.$
  • (A4)
    The function
    $(s,a)\to\int_{S}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)$
    is continuous on $\mathbb{K}$ and there is a constant $c\geq 1$ such that
    $\int_{S}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\leq c\omega(s)^{\theta}$
    for every $(s,a)\in{\mathbb{K}}$ and $c^{\frac{1}{\theta}}\beta<1.$
Case 1: $10<\rho<\gamma<$ 1 .
Note that $10<\theta<$ 1 and for any function $V\in{\cal U}_{+}^{\omega}$, define the operator $F_{1}$ as follows:
$F_{1}V(s):=\max_{a\in A(s)}H_{1}(s,a,V),\ \ s\in S,$
where
$H_{1}(s,a,V):=\left[r(s,a)+\beta\left(\int_{S}V(s^{\prime})q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]^{\theta},\ \ (s,a)\in\mathbb{K}.$ (5)
We can now state our first main result.
Theorem 4
Consider Case 1. Under assumptions (A1)-(A3), the following assertions hold.
$(a)$ The Bellman equation (3) has a unique solution $v_{*}\in{\cal U}_{+}^{\omega}$.
$(b)$ The mapping $F_{1}:{\cal U}_{+}^{\omega}\to{\cal U}_{+}^{\omega}$ is a $c\beta^{\theta}$-contraction. It has a unique fixed point $w_{*}$ and $v_{*}=w_{*}^{\frac{1}{1-\gamma}}$. Moreover, if $F^{n}_{1}$ is the $n$-th composition of $F_{1}$ with itself, then
$\|F_{1}^{n}\hat{0}-w_{*}\|_{\omega}\leq\frac{c^{n}\beta^{n\theta}}{1-c\beta^{\theta}}M.$ (6)
$(c)$ There exists $f_{*}\in\Phi$ that satisfies the optimality equation
$v_{*}(s)=H(s,f_{*}(s),v_{*})\quad\mbox{for all}\quad s\in S.$
$(d)$ The stationary policy $f_{*}$ is optimal.
Case 2: $10<\gamma<\rho<$ 1 .
Note that $\theta>1$ and for any function $V\in{\cal U}_{+}^{\omega}$, define the operator $F_{2}$ as follows:
$F_{2}V(s):=\max_{a\in A(s)}H_{2}(s,a,V),\ \ s\in S,$
where
$H_{2}(s,a,V):=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}V(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right],\ \ (s,a)\in\mathbb{K}.$ (7)
We assume the integrability of the function $V(\cdot)^{\theta}$ with respect to $q(\cdot|s,a)$ for $(s,a)\in\mathbb{K}.$
We can now state our second main result.
Theorem 5
Consider Case 2. Under assumptions (A1)-(A2) and (A4), the following assertions hold.
$(a)$ The Bellman equation (3) has a unique solution $v_{*}\in{\cal U}_{+}^{\omega}$.
$(b)$ The mapping $F_{2}:{\cal U}_{+}^{\omega}\to{\cal U}_{+}^{\omega}$ is a $c^{\frac{1}{\theta}}\beta$-contraction. It has a unique fixed point $w_{*}$ and $v_{*}=w_{*}^{\frac{1}{1-\rho}}$. Moreover, if $F^{n}_{2}$ is the $n$-th composition of $F_{2}$ with itself, then
$\|F_{2}^{n}\hat{0}-w_{*}\|_{\omega}\leq\frac{c^{\frac{n}{\theta}}\beta^{n}}{1-c^{\frac{1}{\theta}}\beta}M.$ (8)
$(c)$ There exists $f_{*}\in\Phi$ that satisfies the optimality equation
$v_{*}(s)=H(s,f_{*}(s),v_{*})\quad\mbox{for all}\quad s\in S.$
$(d)$ The stationary policy $f_{*}$ is optimal.
If $\gamma,\ \rho>1,$ then we must make some changes in our assumptions.
  • (B1)
    For each $s\in S$, the set $A(s)$ is compact and, moreover, the correspondence $s\to A(s)$ is continuous.
  • (B2)
    There exists a continuous function $\omega:S\to[1,\infty)$ such that
    $r(s,a)\leq M\omega(s),\ \ \mbox{for all}\ \ (s,a)\in\mathbb{K}$
    and for some constant $M>0.$ The function $r$ is positive and continuous on $\mathbb{K}.$
  • (B3)
    The function
    $(s,a)\to\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)$
    is continuous and there is a constant $c\geq 1$ such that
    $\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)\leq c\omega(s)$
    for all $(s,a)\in{\mathbb{K}}$ and $c\beta^{\theta}<1.$
  • (B4)
    The function
    $(s,a)\to\int_{S}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)$
    is continuous on $\mathbb{K}$ and there is a constant $c\geq 1$ such that
    $\int_{S}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\leq c\omega(s)^{\theta}$
    for every $(s,a)\in{\mathbb{K}}$ and $c^{\frac{1}{\theta}}\beta<1.$
The difference between (A3) and (B3) and between (A4) and (B4) lies in the properties of the set $\mathbb{K}.$
Case 3: $\gamma>\rho>1.$
Now we have $\theta>1.$ Therefore, the Bellman equation in (3) has the following equivalent form:
$w(s)=\min_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}w(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]$ (9)
for all $s\in S.$ For any function $V\in{\cal C}_{+}^{\omega}$, define the operator $F_{3}$ as follows:
$F_{3}V(s):=\min_{a\in A(s)}H_{3}(s,a,V),\ \ s\in S,$
where
$H_{3}(s,a,V):=r(s,a)+\beta\left(\int_{S}V(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}$ (10)
for all $(s,a)\in\mathbb{K}.$ We shall show that $F_{3}$ maps ${\cal C}_{+}^{\omega}$ into itself. Our next result is as follows.
Theorem 6
Consider Case 3. Under assumptions (B1)-(B2) and (B4), the following assertions hold.
$(a)$ The Bellman equation (3) has a unique positive solution $v_{*}\in{\cal C}_{+}^{\omega}$.
$(b)$ The mapping $F_{3}:{\cal C}_{+}^{\omega}\to{\cal C}_{+}^{\omega}$ is a $c^{\frac{1}{\theta}}\beta$-contraction. It has a unique fixed point $w_{*}$ and $v_{*}=w_{*}^{\frac{1}{1-\rho}}$. Moreover, if $F^{n}_{3}$ is the $n$-th composition of $F_{3}$ with itself, then
$\|F_{2}^{n}\hat{0}-w_{*}\|_{\omega}\leq\frac{c^{\frac{n}{\theta}}\beta^{n}}{1-c\beta}M.$ (11)
$(c)$ There exists $f_{*}\in\Phi$ that satisfies the optimality equation
$v_{*}(s)=H(s,f_{*}(s),v_{*})\quad\mbox{for all}\quad s\in S.$
$(d)$ The stationary policy $f_{*}$ is optimal.
Case 4: $\rho>\gamma>1.$
Now we have $\theta\in(0,1),$ but $01-\gamma<$ 0 and $01-\rho<$ 0 . Define the operator $F_{4}:{\cal C}_{+}^{\omega}\to{\cal C}_{+}^{\omega}$ as follows:
$F_{4}V(s):=\min_{a\in A(s)}H_{4}(s,a,V),\ \ s\in S,$
where
$H_{4}(s,a,V):=\left[r(s,a)+\beta\left(\int_{S}V(s^{\prime})q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]^{\theta},\ \ (s,a)\in\mathbb{K}.$ (12)
Now we state our last main result in this section.
Theorem 7
Consider Case 4. Under assumptions (B1)-(B3), the following assertions hold.
$(a)$ The Bellman equation (3) has a unique solution $v_{*}\in{\cal C}_{+}^{\omega}.$
$(b)$ The mapping $F_{4}:{\cal C}_{+}^{\omega}\to{\cal C}_{+}^{\omega}$ is a $c\beta^{\theta}$-contraction. It has a unique fixed point $w_{*}$ and $v_{*}=w_{*}^{\frac{1}{1-\gamma}}$. Moreover, if $F^{n}_{4}$ is the $n$-th composition of $F_{4}$ with itself, then
$\|F_{4}^{n}\hat{0}-w_{*}\|_{\omega}\leq\frac{c^{n}\beta^{n\theta}}{1-c\beta^{\theta}}M.$ (13)
$(c)$ There exists $f_{*}\in\Phi$ that satisfies the optimality equation
$v_{*}(s)=H(s,f_{*}(s),v_{*})\quad\mbox{for all}\quad s\in S.$
$(d)$ The stationary policy $f_{*}$ is optimal.
The proofs of Theorems 4-7 are given in Section 3.
Remark 8
The assumptions imposed in our work with the weight function $\omega$ have been extensively used in the literature, see Wessels (1977), Durán (2000) and Jaśkiewicz and Nowak (2011). If the function $r$ is bounded, then we can take $\omega(x)\equiv 1.$
Remark 9
From (B2) it follows that
$r_{0}(s):=\min_{a\in A(s)}r(s,a)>0\ \ \mbox{ for all}\ \ s\in S.$
Therefore, the operators $F_{3}$ and $F_{4}$ are considered on ${\cal C}_{+}^{\omega}.$ In particular, $F_{3}\hat{0}$ and $F_{4}\hat{0}$ are well-defined. Therefore, the fixed point of $F_{3}$ or $F_{4}$ can be obtained by iterating these operators on the null function, which makes the approximation procedure simpler.
Remark 10
Cases 1 and 3 were considered in Ren and Stachurski (2020, 2021). Their proof of the existence of a unique solution to the Bellman equation is based on the fixed point theorem of Du (1990) for monotone convex or concave operators on ordered Banach spaces. They also show the geometric convergence of iterations of the corresponding operators. However, this speed of convergence is slower than in Theorems 4 and 6. We discuss this issue in Subsection 4.2. The results of Ren and Stachurski (2020, 2021) also apply to the case when $0<\gamma<1<\rho$ which is not covered by our approach. On the other hand, Cases 2 and 4 cannot be solved by using the theorem of Du (1990), since the corresponding operators are in these aforementioned cases neither concave nor convex. For more comments on this issue see Subsection 4.1.
Remark 11
Note that our conditions in Theorem 6 are weaker than those in Ren and Stachurski (2020), who assumed that the reward function is bounded from below by some function greater than 1. Moreover, they also impose that the integral of the weight function is bounded from below by this weight function multiplied by some constant smaller than $1/\beta^{\theta}.$ The reader is referred to Assumption 4.1 and inequalities (42) and (44) in Ren and Stachurski (2020).
Remark 12
In the most empirically relevant parametric specifications $\rho<\gamma$, implying that the agent prefers early resolution of uncertainty (Cases 1 and 3). For a discussion of this topic the reader is referred to Bansal and Yaron (2004) or Schorfheide et al. (2018); Skiadas (2009). Cases 3 and 4 where the coefficient of relative risk aversion $\gamma>1$ and the intertemporal elasticity of substitution $\frac{1}{\rho}<1$ are also relevant, see Bansal and Yaron (2004); Basu and Bundick (2017); Farhi and Werning (2008).
2.4 An extension to models with non-stationary policies
A definition of the decision process involving the Epstein-Zin aggregators in a framework of general class of history dependent policies is not obvious. For some specifications of the parameters $\gamma$ and $\rho$, one can define utilities in finite step models and then consider the limit as horizon tends to infinity. Then using iterations of some appropriately defined dynamic programming operators, one can prove that an optimal stationary policy derived from the Bellman equation is also optimal in the class of all policies.
A similar construction in risk-sensitive Markov decision processes involving the additive (affine) aggregator and entropic risk measure as a certainty equivalent is described in Asienkiewicz and Jaśkiewicz (2017); Bäuerle and Jaśkiewicz (2018). Additional comments on this issue can be found in Sargent and Stachurski (2024).
A policy is a sequence of Borel measurable mappings $\pi=(\pi_{n})_{n\in\mathbb{N}}$ from the history space to the action set. More precisely, each $\pi_{n}(h_{n})\in A(s_{n})$, $n\in\mathbb{N},$ where $h_{n}=(s_{1},a_{1},\ldots,s_{n-1},a_{n-1},s_{n})$ is the history of the process up to the $n$-th state. For $n=1,$ $h_{1}=s_{1}.$ The class of all policies is denoted by $\Pi$. A Markov policy is a sequence $\pi=(\pi_{n})_{n\in\mathbb{N}}$ where each $\pi_{n}$ depends only on the $n$-th state $s_{n}\in S.$ The class of all Markov policies is denoted by $\Pi_{M}$.
To avoid more complex notation, we shall confine ourselves to consideration of Markov policies. The proof that Markov policies are enough in dynamic programming is standard and will be omitted in this paper. With respect to this issue consult with Bertsekas (2022); Hernández-Lerma and Lasserre (1996); Sargent and Stachurski (2024).
Let us fix a Markov policy $\pi=(\pi_{n})_{n\in\mathbb{N}}$ and define the operator
${\cal T}_{\pi_{n}}\psi(s_{n}):=\left[r(s_{n},\pi_{n}(s_{n}))+\beta\left(\int_{S}\psi(s_{n+1})^{1-\gamma}q(ds_{n+1}|s_{n},\pi_{n}(s_{n}))\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}}$ (14)
for all $s_{n}\in S.$ We assume again the integrability of the function $\psi(\cdot)^{1-\gamma}$ with respect to $q(\cdot|s,a)$ for every $(s,a)\in\mathbb{K}.$ If $1-\gamma<0,$ then $\psi$ must be positive.
The operator in (14) is monotone for some values of parameters $\gamma$ and $\rho$. Here, we shall consider cases discussed in the previous section.
Case (A): $\gamma,\ \rho\in(0,1).$ Moreover, (A1)-(A2) are satisfied and if $\theta\in(0,1)$, then (A3) holds and if $\theta>1,$ then (A4) holds.
Let $\pi=(\pi_{n})_{n\in\mathbb{N}}\in\Pi_{M}$ and $n\geq 1.$ Then the recursive utility in the finite time horizon is defined by the following composition of the operators
$U_{n}(\pi)(s_{1}):={\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\hat{0}(s_{1}).$ (15)
Here, the process starts at state $s_{1}\in S$ and terminates after $n$ steps. with the terminal utility which equals zero. Since $\gamma,\ \rho\in(0,1),$ it follows that $U_{m}(\pi)(s_{1})\leq U_{m+1}(\pi)(s_{1})$ for all $m\geq 1.$ The lifetime recursive utility of an agent is defined as
$U(\pi)(s_{1}):=\lim_{m\to\infty}U_{m}(\pi)(s_{1}),\ \ s_{1}\in S,\ \pi\in\Pi_{M}.$
Definition 13
A policy $\pi_{*}$ is optimal, if
$U(\pi_{*})(s_{1})\geq U(\pi)(s_{1})\quad\mbox{for all}\quad\pi\in\Pi,\ s_{1}\in S.$
The composition ${\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\psi(s_{1})$ can be defined for any $\psi\geq 0$ and we have
${\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\psi(s_{1})\geq{\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\hat{0}(s_{1}).$ (16)
[…]
$U(f)(s_{1})=H(s,f(s_{1}),U(f))\ \ \mbox{for all }\ \ s_{1}\in S.$ (17)
Let us define
${\cal T}\psi(s_{n}):=\max_{a_{n}\in A(s_{n})}H(s_{n},a_{n},\psi),\ \ \psi\geq 0.$ (18)
Note that the sequence ${\cal T}^{m}\hat{0}$ of $m$ compositions of the operator ${\cal T}$ with itself is non-decreasing and for any Markov policy $\pi$, we have
$U(\pi)(s_{1})=\lim_{m\to\infty}{\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{m}}\hat{0}(s_{1})\leq\lim_{m\to\infty}{\cal T}^{m}\hat{0}(s_{1}),\ \ \mbox{for all}\ \ s_{1}\in S.$ (19)
Theorem 14
In Case (A), there exists a stationary optimal policy in the class of all policies.
Proof  As already mentioned, it is enough to prove optimality in the class of all Markov policies. By Theorems 4 and 5, the Bellman equation has a unique solution $v_{*}$ and there exists some $f_{*}\in\Phi$ such that
$v_{*}(s_{1})={\cal T}v_{*}(s_{1})=H(s_{1},f_{*}(s_{1}),v_{*})\quad\mbox{for all}\quad s_{1}\in S.$
From the previous proofs of uniqueness of the solution to the equation
$v(s_{1})=H(s_{1},f_{*}(s_{1}),v)$
we know that $v(s)=U(f_{*})(s)$ for all $s\in S.$ Let $\pi=(\pi_{n})_{n\in\mathbb{N}}$ be any Markov policy. We have $v_{*}(s)={\cal T}v_{*}(s)\geq{\cal T}_{\pi_{n}}v_{*}(s)$ for all $n\in\mathbb{N}$ and $s\in S.$ Iterating these inequalities we obtain for any $m\in\mathbb{N}$ that
$v_{*}(s)$$\geq$ ${\cal T}_{\pi_{1}}v_{*}(s_{1})\geq{\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}v_{*}(s_{1})\geq\cdots\geq{\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{m}}v_{*}(s_{1})$
$\geq$ ${\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{m}}\hat{0}(s_{1})=U_{m}(\pi)(s_{1}).$
Hence, it follows that
$U(f_{*})(s_{1})=v_{*}(s_{1})\geq\lim_{m\to\infty}U_{m}(\pi)(s_{1})=U(\pi)(s_{1})$
for every $s_{1}\in S.$Box
Case (B): $\gamma>1$ and $\rho>1.$ Moreover, (B1)-(B2) are satisfied and if $\theta\in(0,1),$ then (B3) holds and if $\theta>1,$ then (B4) holds.
For any $f\in\Phi$ we define ${\cal T}_{f}\hat{0}(s):=r(s,f(s))^{\frac{1}{1-\rho}}$ for $s\in S.$ For $\psi>0$ we put ${\cal T}_{f}\psi(s):=H(s,f(s),\psi).$ Then for any Markov policy $\pi=(\pi_{n})_{n\in\mathbb{N}}\in\Pi_{M}$, the recursive utility $U_{n}(\pi)(s_{1})$ can be defined as in (15). It should be noted that if $01-\gamma<$ 0 and $1-\rho<0,$ then the sequence
$U_{n}(\pi)(s_{1})={\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\hat{0}(s_{1}),\ \ n\in\mathbb{N},$
is non-increasing, and therefore, we can define
$U(\pi)(s_{1}):=\lim_{m\to\infty}U_{m}(\pi)(s_{1}).$
The sequence ${\cal T}^{m}\hat{0}$ of $m$ iterations of the operator ${\cal T}$ defined in (18) with itself is non-increasing. Under our assumptions ${\cal T}^{n}\hat{0}(\cdot)$ is continuous on $S$ for every $n\in\mathbb{N}.$ Let
$w_{0}(s):=\lim_{m\to\infty}{\cal T}^{m}\hat{0}(s),\quad s\in S.$
The function $w_{0}$ is upper semicontinuous and using Proposition 10.1 in Schäl (1975), we infer that
$w_{0}(s)={\cal T}w_{0}(s)=\max_{a\in A(s)}H(s,a,w_{0}),\ \ \mbox{for all}\ \ s\in S.$ (20)
Theorem 15
In Case (B), there exists a stationary optimal policy in the class of all policies.
Proof We can restrict ourselves to Markov policies. By Theorems 6 and 7 the Bellman equation has a unique solution $v_{*}$ and there exists some $f_{*}\in\Phi$ such that
$v_{*}(s_{1})={\cal T}v_{*}(s_{1})=H(s_{1},f_{*}(s_{1}),v_{*})\quad\mbox{for all}\quad s_{1}\in S.$
Our arguments used in the proofs of Theorems 6 and 7 and (20) imply that $w_{0}(s)=v_{*}(s)=U(f_{*})(s)$ for all $s\in S.$ Fix any state $s_{1}\in S$, a number $\epsilon>0$ and a Markov policy $\pi=(\pi_{n})_{n\in\mathbb{N}}.$ There exists $N>0$ such that
$w_{0}(s_{1})=v_{*}(s_{1})\geq{\cal T}^{m}\hat{0}(s_{1})-\epsilon\geq U_{m}(\pi)(s_{1})-\epsilon,\quad\mbox{for all}\quad m>N.$
Hence
$U(f_{*})(s_{1})=v_{*}(s_{1})\geq\lim_{m\to\infty}(U_{m}(\pi)(s_{1})-\epsilon)=U(\pi)(s_{1})-\epsilon.$
Since $\epsilon>0$ can be arbitrarily small, it follows that $U(f_{*})(s_{1})\geq U(\pi)(s_{1}).$Box
3 Proofs of Theorems 4-7
The proof of Theorem 4 is proceeded by an auxiliary lemma.
Lemma 16
Under assumptions (A1)-(A3), the assertions $(a)$-$(b)$ of Theorem 4 hold.
Proof   First, we must show that $F_{1}$ maps ${\cal U}_{+}^{\omega}$ into itself. Let $V\in{\cal U}_{+}^{\omega}.$ Note that from (A2) and (A3), we get
00 0 $\leq$ $F_{1}V(s)\leq\left[M\omega(s)+\beta\left(\int_{S}\|V\|_{\omega}\omega(s^{\prime})q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]^{\theta}$
$\leq$ $\left[M\omega(s)^{\frac{1}{\theta}}+\beta\left(c\|V\|_{\omega}\omega(s)\right)^{\frac{1}{\theta}}\right]^{\theta}\leq\omega(s)\left[M+\beta c^{\frac{1}{\theta}}\|V\|_{\omega}^{\frac{1}{\theta}}\right]^{\theta},\quad s\in S.$
Additionally, by Lemma 8.5.5 in Hernández-Lerma and Lasserre (1999), it follows that the function
$(s,a)\to\int_{S}V(s^{\prime})q(ds^{\prime}|s,a)$
is upper semicontinuous. Therefore, $(s,a)\to H_{1}(s,a,V)$ is upper semicontinuous. Consequently, by (A1) and Lemma 3.8.3(c) in Hernández-Lerma and Lasserre (1999) the function $F_{1}V$ is upper semicontinuous on $S.$ Next we prove that $F_{1}$ is a $c\beta^{\theta}$-contraction. Let $V_{1},V_{2}\in{\cal U}^{\omega}_{+}$. Consider the norm
$\|\boldsymbol{z}\|_{\frac{1}{\theta}}=\left(|z_{1}|^{\frac{1}{\theta}}+|z_{2}|^{\frac{1}{\theta}}\right)^{\theta}$
for any vector $\boldsymbol{z}=(z_{1},z_{2})\in{\mathbb{R}}^{2}.$ Observe that $\|\cdot\|_{\frac{1}{\theta}}$ is indeed the norm, since $\frac{1}{\theta}>1.$ Now take
$\boldsymbol{x}:=\left(r^{\theta},\beta^{\theta}t_{1}\right),\quad\boldsymbol{y}:=\left(r^{\theta},\beta^{\theta}t_{2}\right),$
where we use the following abbreviations
$r^{\theta}:=r(s,a)^{\theta},\quad t_{i}:=\int_{S}V_{i}(s^{\prime})q(ds^{\prime}|s,a)$
for $(s,a)\in\mathbb{K}.$ We notice that
$\|\boldsymbol{x}\|_{\frac{1}{\theta}}=H_{1}(s,a,V_{1})\ \mbox{ and }\ \|\boldsymbol{y}\|_{\frac{1}{\theta}}=H_{1}(s,a,V_{2})$
with $H_{1}$ in (5). Although for convenience of notation we do not indicate dependence of $\boldsymbol{x}$ and $\boldsymbol{y}$ on $(s,a)$ we keep it in mind. Using these facts, we obtain
$|F_{1}V_{1}(s)-F_{1}V_{2}(s)|$$\leq$ $\max_{a\in A(s)}|H_{1}(s,a,V_{1})-H_{1}(s,a,V_{2})|=\max_{a\in A(s)}\left|\|\boldsymbol{x}\|_{\frac{1}{\theta}}-\|\boldsymbol{y}\|_{\frac{1}{\theta}}\right|$
$\leq$ $\max_{a\in A(s)}\|\boldsymbol{x}-\boldsymbol{y}\|_{\frac{1}{\theta}}=\max_{a\in A(s)}\|(0,\beta^{\theta}(t_{1}-t_{2}))\|_{\frac{1}{\theta}}$
$=$ $\max_{a\in A(s)}\left(0+\left|\beta^{\theta}(t_{1}-t_{2})\right|^{\frac{1}{\theta}}\right)^{\theta}=\max_{a\in A(s)}\beta^{\theta}|t_{1}-t_{2}|$
$\leq$ $\max_{a\in A(s)}\beta^{\theta}\int_{S}\frac{|V_{1}(s^{\prime})-V_{2}(s^{\prime})|}{\omega(s^{\prime})}\omega(s^{\prime})q(ds^{\prime}|s,a)$
$\leq$ $\beta^{\theta}c\|V_{1}-V_{2}\|_{\omega}\omega(s),\quad s\in S.$
Now dividing all sides of these inequalities by $\omega(s)$ and taking the supremum over all $s\in S$ we get
$\|F_{1}V_{1}-F_{1}V_{2}\|_{\omega}\leq c\beta^{\theta}\|V_{1}-V_{2}\|_{\omega}.$
Hence, $F_{1}$ is a $c\beta^{\theta}$-contraction. Since ${\cal U}_{+}^{\omega}$ is a closed subset of the Banach space ${\cal C}^{\omega}$ we conclude from the Banach contraction principle that there a unique $w_{*}\in{\cal U}^{\omega}_{+}$ such that
$w_{*}(s)=F_{1}w_{*}(s)\quad s\in S.$
Putting $v_{*}:=w_{*}^{\frac{1}{1-\gamma}}$ we get that $v_{*}$ is non-negative, upper semicontinuous and
$v_{*}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}v_{*}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}}$ (21)
for all $s\in S.$ To finish this proof note that (6) now follows from (1) with $F=F_{1}.$Box
Proof of Theorem 4  In view of Lemma 16 it remains to prove assertions $(c)$ and $(d).$ The existence of a maximiser $f_{*}\in\Phi$ of the right-hand side of (21) follows from our assumptions (A1)-(A3) and Corollary 1 in Brown and Purves (1973). Let ${\cal B}^{\omega}$ be the set of all Borel measurable functions on $S$ for which $\|\cdot\|_{\omega}$-norm is finite and ${\cal B}_{+}^{\omega}$ be the set of non-negative functions in ${\cal B}^{\omega}.$ For any $f\in\Phi$ and $w\in{\cal B}_{+}^{\omega}$, we define the operator $F_{1f}$ as follows
$F_{1f}w(s):=H_{1}(s,f(s),w),\quad s\in S.$
Using the measurability results of Brown and Purves (1973), similarly as above, we can show that $F_{1f}w\in{\cal B}_{+}^{\omega}.$ Moreover, $F_{1f}$ is a $c\beta^{\theta}$-contraction. Therefore, the operator $F_{1f}$ has s a unique fixed point $w_{f}\in{\cal B}_{+}^{\omega}$ and for any function $w_{0}\in{\cal B}_{+}^{\omega}$
$\lim_{n\to\infty}\|F^{n}_{1f}w_{0}-w_{f}\|_{\omega}=0.$
Observe also that $F_{1f}$ is monotone, i.e. $F_{1f}w_{1}\geq F_{1f}w_{2}$ whenever $w_{1}\geq w_{2}$ and $w_{1},w_{2}\in{\cal B}_{+}^{\omega}.$ For any $f\in\Phi$, we have $F_{1f}w_{*}\leq w_{*}.$ By iteration of this inequality we get $F^{n}_{1f}w_{*}\leq w_{*}$ and consequently $w_{f}\leq w_{*},$ since $w_{*}\in{\cal U}_{+}^{\omega}\subset{\cal B}_{+}^{\omega}.$ On the other hand, due to Lemma 16 we have $F_{1f_{*}}w_{*}=w_{*}=F_{1}w_{*}.$ Again iterating $F_{1f_{*}}$ we obtain $\lim_{n\to\infty}F^{n}_{1f_{*}}w_{*}=w_{f_{*}}=w_{*}.$ Summing up,
$w_{f}\leq w_{*}=w_{f_{*}}\quad\mbox{for all }\ f\in\Phi,$
which implies via transformation $v:=w^{\frac{1}{1-\gamma}}$ with $01-\gamma>$ 0 that
$v_{f}\leq v_{*}=v_{f_{*}}\quad\mbox{for all }\ f\in\Phi.$
This means that $f_{*}$ is an optimal stationary policy. $\Box$
The proof of Theorem 5 is proceeded by an auxiliary fact.
Lemma 17
Under assumptions (A1)-(A2) and (A4), the assertions $(a)$-$(b)$ of Theorem 5 hold.
Proof   First, we must show that $F_{2}:{\cal U}_{+}^{\omega}\to{\cal U}_{+}^{\omega}.$ Indeed, note that from (A2) and (A4), we get
$0\leq|F_{2}V(s)|$$\leq$ $\left[M\omega(s)+\beta\left(\int_{S}\|V\|^{\theta}_{\omega}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]$
$\leq$ $\left[M\omega(s)+\beta\left(c\|V\|^{\theta}_{\omega}\omega(s)^{\theta}\right)^{\frac{1}{\theta}}\right]\leq\omega(s)\left[M+\beta c^{\frac{1}{\theta}}\|V\|_{\omega}\right].$
Again, by Lemma 8.5.5 in Hernández-Lerma and Lasserre (1999), it follows that the function
$(s,a)\to\int_{S}V(s^{\prime})q(ds^{\prime}|s,a)$
is upper semicontinuous. Therefore, $(s,a)\to H_{2}(s,a,V)$ is upper semicontinuous. Consequently, by (A1) and the Maximum Theorem of Berge (see Section 17.5 in Aliprantis and Border (2006)), the function $F_{2}V$ is upper semicontinuous on $S.$ Now we prove that $F_{2}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Let $V_{1},V_{2}\in{\cal U}^{\omega}_{+}$. Then, by the Minkowski inequality we have
$|F_{2}V_{1}(s)-F_{2}V_{2}(s)|$$\leq$ $\max_{a\in A(s)}\beta\left|\left(\int_{S}V_{1}(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}-\left(\int_{S}V_{2}(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right|$
$\leq$ $\max_{a\in A(s)}\beta\left(\int_{S}|V_{1}(s^{\prime})-V_{2}(s^{\prime})|^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}$
$\leq$ $\max_{a\in A(s)}\beta\left(\int_{S}\|V_{1}-V_{2}\|_{\omega}^{\theta}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}$
$\leq$ $\beta\|V_{1}-V_{2}\|_{\omega}c^{\frac{1}{\theta}}\omega(s),\quad s\in S.$
Now dividing all sides by $\omega(s)$ and taking the supremum over all $s\in S$ we get
$\|F_{2}V_{1}-F_{2}V_{2}\|_{\omega}\leq c^{\frac{1}{\theta}}\beta\|V_{1}-V_{2}\|_{\omega}.$
Hence, $F$ is $c^{\frac{1}{\theta}}\beta$-contraction. Since ${\cal U}_{+}^{\omega}$ is a closed subset of the Banach space ${\cal C}^{\omega}$ we conclude from the Banach contraction principle that there exists $w_{*}\in{\cal U}^{\omega}_{+}$ such that
$w_{*}(s)=F_{2}w_{*}(s)\quad s\in S.$
Putting $v_{*}:=w_{*}^{\frac{1}{1-\rho}}$ we get that
$v_{*}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}v_{*}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},\quad s\in S.$ (22)
for all $s\in S.$ Note that (8) follows from (1) with $F=F_{2}.$Box
Proof of Theorem 5  In view of Lemma 17 it remains to prove assertions $(c)$ and $(d).$ The existence of a maximiser $f_{*}\in\Phi$ of the right-hand side of (22) follows from our assumptions (A1)-(A2), (A4) and a measurable selection theorem, see Corollary 1 in Brown and Purves (1973). Again we consider the space ${\cal B}_{+}^{\omega}.$ For any $f\in\Phi$ and $w\in{\cal B}_{+}^{\omega}$ we define the operator $F_{2f}$ as follows
$F_{2f}w(s):=H_{2}(s,f(s),w),\quad s\in S.$
Similarly as in the proof of Theorem 4 we may show that $F_{2f}w\in{\cal B}_{+}^{\omega}.$ Moreover, $F_{2f}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Therefore, the operator $F_{2f}$ has s a unique fixed point $w_{f}\in{\cal B}_{+}^{\omega}$ and for any function $w_{0}\in{\cal B}_{+}^{\omega}$
$\lim_{n\to\infty}\|F^{n}_{2f}w_{0}-w_{f}\|_{\omega}=0.$
Observe also that $F_{2f}$ is monotone, i.e. $F_{2f}w_{1}\geq F_{2f}w_{2}$ whenever $w_{1}\geq w_{2}$ and $w_{1},w_{2}\in{\cal B}_{+}^{\omega}.$ For any $f\in F$ we have $F_{2f}w_{*}\leq w_{*}.$ By iteration of this inequality we get $F^{n}_{2f}w_{*}\leq w_{*}$ and consequently $w_{f}\leq w_{*}.$ On the other hand, $F_{2f_{*}}w_{*}=w_{*}=F_{2}w_{*}.$ Again iterating $F_{2f_{*}}$ we obtain $\lim_{n\to\infty}F^{n}_{2f_{*}}w_{*}=w_{f_{*}}=w_{*}.$ Thus,
$w_{f}\leq w_{*}=w_{f_{*}}\quad\mbox{for all }\ f\in\Phi,$
which via transformation $v:=w^{\frac{1}{1-\rho}}$ with $01-\rho>$ 0 gives that
$v_{f}\leq v_{*}=v_{f_{*}}\quad\mbox{for all }\ f\in\Phi.$
This means that $f_{*}$ is an optimal stationary policy. $\Box$
Proof of Theorem 6  Proceeding as in Lemma 17 we obtain that $F_{3}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Moreover, from (B2) and (B4) for any $V\in{\cal C}_{+}^{\omega}$ we have
$0<\|F_{3}V\|_{\omega}\leq M+\beta c^{\frac{1}{\theta}}\|V\|_{\omega}.$
Note that Lemma 8.5.5 in Hernández-Lerma and Lasserre (1999), (B1)-(B2), (B4) and the Maximum Theorem of Berge (Section 17.5 in Aliprantis and Border (2006)) imply that $F_{3}V$ is continuous on $S.$ Since ${\cal C}_{+}^{\omega}$ is a closed subset of the Banach space ${\cal C}^{\omega}$ we conclude from the Banach contraction principle that there exists $w_{*}\in{\cal C}^{\omega}_{+}$ such that
$w_{*}(s)=F_{3}w_{*}(s)\quad s\in S.$
Note that $w_{*}(s)\geq\min_{a\in A(s)}r(s,a)>0$ for every $s\in S.$ Putting $v_{*}:=w_{*}^{\frac{1}{1-\rho}}$ we get that
$v_{*}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}v_{*}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},\quad s\in S.$ (23)
for all $s\in S.$ Clearly, (11) follows from (1) with $F=F_{3}.$ Now, we prove (c) and (d). Let $f_{*}\in\Phi$ be a maximiser of the right-hand side of (23). For any $f\in\Phi$ and $w\in{\cal B}_{+}^{\omega}$ set
$F_{3f}w(s):=H_{3}(s,f(s),w),\quad s\in S.$
It follows that $F_{3f}w\in{\cal B}_{+}^{\omega}$ and $F_{3f}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Therefore, the operator $F_{3f}$ has s a unique fixed point $w_{f}\in{\cal B}_{+}^{\omega}$ and for any function $w_{0}\in{\cal B}_{+}^{\omega}$
$\lim_{n\to\infty}\|F^{n}_{3f}w_{0}-w_{f}\|_{\omega}=0.$
Additionally, $F_{3f}$ is also monotone, i.e. $F_{3f}w_{1}\geq F_{3f}w_{2}$ if $w_{1}\geq w_{2}$ and $w_{1},w_{2}\in{\cal B}_{+}^{\omega}.$ For any $f\in F$ we have $F_{3f}w_{*}\geq w_{*}.$ By iteration of this inequality we get $F^{n}_{3f}w_{*}\geq w_{*}$ and consequently $w_{f}\geq w_{*}.$ On the other hand, $F_{3f_{*}}w_{*}=w_{*}=F_{3}w_{*}.$ Again iterating $F_{3f_{*}}$ we obtain $\lim_{n\to\infty}F^{n}_{3f_{*}}w_{*}=w_{f_{*}}=w_{*}.$ Thus,
$w_{f}\geq w_{*}=w_{f_{*}}\quad\mbox{for all }\ f\in\Phi,$
which via transformation $v:=w^{\frac{1}{1-\rho}}$ with $01-\rho<$ 0 yields that
$v_{f}\leq v_{*}=v_{f_{*}}\quad\mbox{for all }\ f\in\Phi.$
This means that $f_{*}$ is an optimal stationary policy. $\Box$
Proof of Theorem 7  Similarly as in Lemma 16 we show that $F_{4}$ is $c\beta^{\theta}$-contraction. Moreover, from (B2) and (B3) for any $V\in{\cal C}_{+}^{\omega}$ we have
$0<\|F_{4}V\|_{\omega}\leq\left[M+\beta c^{\frac{1}{\theta}}\|V\|_{\omega}^{\frac{1}{\theta}}\right]^{\theta}.$
Lemma 8.5.5 in Hernández-Lerma and Lasserre (1999), (B1)-(B3) and the Maximum Theorem of Berge (Section 17.5 in Aliprantis and Border (2006)) imply the continuity of $F_{4}V$ on $S.$ From the Banach contraction principle there exists $w_{*}\in{\cal C}^{\omega}_{+}$ such that
$w_{*}(s)=F_{4}w_{*}(s)\quad s\in S.$
By (B2) $w_{*}(s)\geq\min_{a\in A(s)}r(s,a)>0$ for every $s\in S.$ Putting $v_{*}:=w_{*}^{\frac{1}{1-\gamma}}$ we get that
$v_{*}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}v_{*}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},\quad s\in S.$ (24)
for all $s\in S.$ Obviously, (13) follows from (1) with $F=F_{4}.$ It remains to prove (c) and (d). Let $f_{*}\in\Phi$ be a maximiser of the right-hand side of (24). For $f\in\Phi$ and $w\in{\cal B}_{+}^{\omega}$ set
$F_{4f}w(s):=H_{4}(s,f(s),w),\quad s\in S.$
Then, $F_{4f}w\in{\cal B}_{+}^{\omega}$ and $F_{4f}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Hence, the operator $F_{4f}$ has s a unique fixed point $w_{f}\in{\cal B}_{+}^{\omega}$ and for any function $w_{0}\in{\cal B}_{+}^{\omega}$
$\lim_{n\to\infty}\|F^{n}_{4f}w_{0}-w_{f}\|_{\omega}=0.$
Additionally, $F_{4f}$ is also monotone, i.e. $F_{4f}w_{1}\geq F_{4f}w_{2}$ if $w_{1}\geq w_{2}$ and $w_{1},w_{2}\in{\cal B}_{+}^{\omega}.$ For any $f\in F$ we have $F_{4f}w_{*}\geq w_{*}.$ By iteration we obtain $F^{n}_{4f}w_{*}\geq w_{*}$ and finally $w_{f}\geq w_{*}.$ On the other hand, $F_{4f_{*}}w_{*}=w_{*}=F_{4}w_{*}.$ Iterating $F_{4f_{*}}$ we get $\lim_{n\to\infty}F^{n}_{4f_{*}}w_{*}=w_{f_{*}}=w_{*}.$ Thus,
$w_{f}\geq w_{*}=w_{f_{*}}\quad\mbox{for all }\ f\in\Phi,$
which via transformation $v:=w^{\frac{1}{1-\gamma}}$ with $01-\gamma<$ 0 yields that
$v_{f}\leq v_{*}=v_{f_{*}}\quad\mbox{for all }\ f\in\Phi.$
This proves that $f_{*}$ is an optimal stationary policy. $\Box$
4 Remarks on an application of Du’s fixed point theorem and Banach’s contraction principle
In this section, we consider bounded utility functions.
4.1 Du’s theorem and convex/concave operators
[…]
Proposition 18
Let $T:I\to I$ be a monotone increasing operator. If either (i) $T$ is convex on $I$ and for some $\epsilon\in(0,1)$
$Tg_{2}\leq(1-\epsilon)g_{2}+\epsilon g_{1},$ (25)
or (ii) $T$ is concave on $I$ and for some $\epsilon\in(0,1)$
$Tg_{1}\geq(1-\epsilon)g_{1}+\epsilon g_{2},$ (26)
then $T$ has a unique fixed point $v_{*}\in I.$ Moreover, for any $v_{0}\in I$ and the sequence defined by $v_{n}=Tv_{n-1},$ $n\in\mathbb{N},$ there exists $B>0$ such that
$\|v_{n}-v_{*}\|\leq B(1-\epsilon)^{n},\ \ n\in\mathbb{N}.$ (27)
Since ${\cal C}_{+}$ is a normal cone in $\cal C$ with constant $N=1$, from the proofs of the theorem of Du (1990) or Theorem 2.1.2 in Zhang (2012), it follows that, if $T$ satisfies $(i),$ then
$B=\|g_{1}-g_{2}\|+\frac{2\|Tg_{2}-g_{2}\|}{\epsilon^{2}}$ (28)
and if $T$ satisfies $(ii),$ then
$B=\|g_{1}-g_{2}\|+\frac{2\|Tg_{1}-g_{1}\|}{\epsilon^{2}}.$ (29)
Remark 19
Inequalities (25) and (26) are called boundary conditions. They are crucial in proving the geometric convergence of iterations of the operator $T$ to a fixed point that belongs to the interior of the set $I.$
Below we consider two cases discussed in Ren and Stachurski (2020) who applied the transformation $v=\widetilde{v}^{1-\gamma}$. Under this transformation $H$ defined in (2) can be written as
${\cal H}(s,a,v)=\left[r(s,a)+\beta\left(\int_{S}v(s^{\prime})q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]^{\theta},\quad(s,a)\in\mathbb{K}.$
Recall also that $\theta=\frac{1-\gamma}{1-\rho}.$
Case 1: $0<\rho<\gamma<1,$ i.e., $\theta\in(0,1).$ In this setting, Bellman equation (4) is equivalent to
$v(s)=Tv(s):=\max_{a\in A(s)}{\cal H}(s,a,v),$ (30)
From the convexity of the function $\psi(t)=(b+\beta t^{1/\theta})^{\theta}$ on $\mathbb{R}_{+}$ for any $b\geq 0$, it follows that for each $v_{1}$, $v_{2}$ and $\xi\in(0,1)$,
${\cal H}(s,a,\xi v_{1}+(1-\xi)v_{2})\leq\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2}).$ (31)
Hence,
$T(\xi v_{1}(s)+(1-\xi)v_{2})(s)\leq\xi Tv_{1}(s)+(1-\xi)Tv_{2}(s),$
i.e., $T$ is convex.
Case 3: $1<\rho<\gamma,$ i.e., $\theta>1.$ In this setting, Bellman equation (4) is equivalent to
$v(s)=Tv(s):=\min_{a\in A(s)}{\cal H}(s,a,v),\quad s\in S.$ (32)
Since the function $\psi(t)=(b+\beta t^{1/\theta})^{\theta}$ is concave on $\mathbb{R}_{+}$ for any $b\geq 0,$ it follows that for each $v_{1}$, $v_{2}$ and $\xi\in(0,1)$,
${\cal H}(s,a,\xi v_{1}+(1-\xi)v_{2})\geq\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2}).$ (33)
Hence,
$T(\xi v_{1}(s)+(1-\xi)v_{2})(s)\geq\xi Tv_{1}(s)+(1-\xi)Tv_{2}(s),$
i.e., $T$ is concave.
Below we comment on two other cases for which the Banach contraction mapping theorem is applied in Subsection 2.3 after suitable transformations. We explain why the fixed point theorem of Du (1990) cannot be applied.
Case 2: $0<\gamma<\rho<1,$ i.e., $\theta>1.$ Now Bellman equation (4) is equivalent to (30). Now $v\to{\cal H}(s,a,v)$ is concave. Therefore, (33) holds for each $v_{1}$, $v_{2}$ and $\xi\in(0,1).$ Hence,
$\max_{a\in A(s)}{\cal H}(s,a,\xi v_{1}+(1-\xi)v_{2})\geq\max_{a\in A(s)}(\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2})),$
but
$\max_{a\in A(s)}(\xi{\cal H}(s,a,v_{1})+(1-\xi{\cal H}(s,a,v_{2}))\not\geq\xi\max_{a\in A(s)}{\cal H}(s,a,v_{1})+(1-\xi)\max_{a\in A(s)}{\cal H}(s,a,v_{2}).$
Thus, in this case $T$ is not concave and the theorem of Du (1990) cannot be applied.
Case 4: $1<\gamma<\rho,$ i.e., $\theta\in(0,1).$ Here, Bellman equation (4) is equivalent to (32). The function $\psi(t)=(b+\beta t^{1/\theta})^{\theta}$ is convex on $\mathbb{R}_{+}$ and therefore, (31) holds for each $v_{1}$, $v_{2}$ and $\xi\in(0,1).$ Hence,
$T(\xi v_{1}(s)+(1-\xi)v_{2})(s)\leq\min_{a\in A(s)}\left(\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2})\right)$
but
$\min_{a\in A(s)}\left(\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2})\right)\not\leq\xi\min_{a\in A(s)}{\cal H}(s,a,v_{1})+(1-\xi)\min_{a\in A(s)}{\cal H}(s,a,v_{2}).$
Thus, $T$ is not convex and the theorem of Du (1990) cannot be applied to this case.
4.2 Comparing the speed of convergence in Du’s theorem and Banach’s theorem
Our aim in this section is to compare the iteration algorithms proposed by Ren and Stachurski (2020, 2021) with our modifications. We are interested in estimating the speed of convergence in both cases.
We remind that from the proof of the Banach contraction mapping principle, it follows that, if $F:{\cal C}_{+}\to{\cal C}_{+}$ is a $\delta$-contraction, then $F$ has a unique fixed point $w_{*}\in{\cal C}_{+}$ and for each $n\in\mathbb{N}$
$\|w_{n}-w_{*}\|\leq L\delta^{n}\quad\mbox{where}\quad L=\frac{\|F\hat{0}-\hat{0}\|}{1-\delta},$ (34)
$w_{0}\equiv hat{0}$ and $w_{n}=Fw_{n-1},$ $n\in\mathbb{N}.$
Case 1: $10<\rho<\gamma<$ 1 . We assume that
$m=\min_{(s,a)\in\mathbb{K}}r(s,a)=0\quad\mbox{and}\quad M=\max_{(s,a)\in\mathbb{K}}r(s,a).$
We fix $y>0$ and similarly as in Ren and Stachurski (2020) we consider the constant functions
$g_{1}(s):=\hat{0}\quad\mbox{and}\quad g_{2}(s):=\left(\frac{M+y}{1-\beta}\right)^{\theta},\ \ s\in S.$
As noted by Ren and Stachurski (2020), the operator $T$ considered in this case is convex. Thus, the geometric convergence of iterations of $T$ with $v_{0}=g_{1}=\hat{0}$ can be analysed using the formula (28). Below we try to estimate the values of $B$ and $1-\epsilon$ when $\theta=1/2.$
Example 1
Assume that $M=1$, $\beta=0.9,$ $\rho=1/2$ and $\gamma=3/4.$ Then $\theta=1/2.$ Let
$\min_{s\in S}\max_{a\in A(s)}r(s,a)=0.$
We have
$\|g_{1}-g_{2}\|=\sqrt{10(1+y)}\quad\mbox{and}\quad Tg_{2}(s)=\sqrt{\max_{a\in A(s)}r(s,a)+9(1+y)}$
and consequently,
$\|Tg_{2}-g_{2}\|$ $=$ $\sqrt{10+10y}-\sqrt{9(1+y)+\min_{s\in S}\max_{a\in A(s)}r(s,a)}$
$=$ $\sqrt{10+10y}-3\sqrt{1+y}=(\sqrt{10}-3)\sqrt{1+y}.$
It is easy to see that (25) holds if
$\sqrt{1+9(1+y)}\leq(1-\epsilon)\sqrt{10+10y}.$
Hence,
$\epsilon\leq\epsilon(y):=1-\sqrt{\frac{1+0.9y}{1+y}},\quad y>0.$
Then
$1-\epsilon(y)=\sqrt{\frac{1+0.9y}{1+y}}.$
To obtain the fastest geometric convergence of iterations of the operator $T$ with starting function $v_{0}=\hat{0}$, we apply the following method. We can minimise $B(y)(1-\epsilon(y))$ where $B(y)=B$ is given in (28). Note that
$B(y)=B=\sqrt{10+10y}+\frac{2(\sqrt{10}-3)\sqrt{1+y}}{\left(1-\sqrt{\frac{1+0.9y}{1+y}}\right)^{2}}$
and
$B(y)(1-\epsilon(y))=\left(\sqrt{10}+\frac{2(\sqrt{10}-3)}{\left(1-\sqrt{\frac{1+0.9y}{1+y}}\right)^{2}}\right)\sqrt{1+0.9y}$ (35)
where $y>0.$ The minimum value of $B(y)(1-\epsilon(y))$ over the interval $(0,\infty)$ is attained at $y_{0}\approx 4.10707$ and equals $B(y_{0})(1-\epsilon(y_{0}))\approx 424.197.$ Next $1-\epsilon(y_{0})\approx 0.958947704>\sqrt{\beta}=\sqrt{0.9}\approx 0.948683298$ The reason for these results is that $T$ is not treated as a contraction mapping in Proposition 18. We note that
$\inf_{y>0}(1-\epsilon(y))=\lim_{y\to\infty}(1-\epsilon(y))=\sqrt{\beta}=\sqrt{0.9},$
but $\lim_{y\to\infty}B(y)=\infty.$ Thus, minimising only the expression $1-\epsilon(y)$ makes no sense.
The operator $F_{1}$ proposed in Subsection 2.3 to study iterations with starting function $v_{0}=\hat{0}$ is as $T$ in (30) with $\theta=1/2.$ We have shown that $F_{1}$ is a $\sqrt{\beta}$-contraction. Since the algorithm using the operator $F_{1}$ starts with $v_{0}=\hat{0},$ the constant $L$ in (34) is equal to 1/($19.48681/(1-\sqrt{\beta})=1/(1-\sqrt{0.9})=$ 19.4868 . It is obvious that the speed of convergence of iterations of the operator $F_{1}$ viewed as a $\sqrt{\beta}$-contraction is much better than in the former case. It should be noted that the solution to Bellman equation (4) in this example is $w_{*}^{4}.$
Case 3: $1<\rho<\gamma.$ We now assume that
$0<m=\min_{(s,a)\in\mathbb{K}}r(s,a)\quad\mbox{and}\quad M=\max_{(s,a)\in\mathbb{K}}r(s,a).$
We fix $x\in(0,m)$ and as in Ren and Stachurski (2020) we consider the constant functions
$g_{1}(s):=\left(\frac{x}{1-\beta}\right)^{\theta}\quad\mbox{and}\quad g_{2}(s):=\left(\frac{M}{1-\beta}\right)^{\theta}.$
As noted by Ren and Stachurski (2020), the operator $T$ considered in this case is concave. Therefore, the geometric convergence of iterations of $T$ with $v_{0}=g_{1}$ can be analysed using the formula (27) with $B$ defined in (29). Below we try to estimate $B$ and $1-\epsilon$ in (29) when $\theta=2.$
Example 2
Assume that $m=1,$ $M=5$ and
$\max_{s\in S}\min_{a\in A(s)}r(s,a)=3.$
Let $\beta=0.9,$ $\rho=1.25$ and $\gamma=1.5.$ Then $\theta=2$ and $x\in(0,1).$ Moreover,
$\|g_{1}-g_{2}\|=100(25-x^{2})\quad\mbox{and}\quad Tg_{1}(s)=\min_{a\in A(s)}(r(s,a)+9x)^{2}\leq(3+9x)^{2}$
for all $s\in S.$ Hence,
$\|Tg_{1}-g_{1}\|=(3+9x)^{2}-100x^{2}=9+54x-19x^{2}.$
One can easily see that (26) holds, if
$(1+9x)^{2}\geq(1-\epsilon)100x^{2}+\epsilon 2500.$
Thus, we have
$\epsilon\leq\epsilon(x):=\frac{1+18x-19x^{2}}{100(25-x^{2})},\quad x\in(0,1),$
and
$1-\epsilon(x)=\frac{2499-18x-81x^{2}}{100(25-x^{2})}.$
To get the best geometric convergence of iterations of the operator $T$ with starting function $v_{0}=g_{1}$, we can apply two approaches. First, we can try to minimise $B(x)(1-\epsilon(x)),$ where $B(x)=B$ is given in (29). We obtain
$B(x)=B=100(25-x^{2})+\frac{2\cdot 100^{2}(25-x^{2})^{2}(9+54x-19x^{2})}{(1+18x-19x^{2})^{2}},$
and
$B(x)(1-\epsilon(x))=(2499-18x-81x^{2})R(x)$
where
$R(x):=\left(1+\frac{200(25-x^{2})(9+54x-19x^{2})}{(1+18x-19x^{2})^{2}}\right),$
for $x\in(0,1).$ The minimum value of $B(x)(1-\epsilon(x))$ over the interval $(0,1)$ is attained at $x_{0}\approx 0.378544$ and is equal to $B(x_{0})(1-\epsilon(x_{0}))\approx 1.27143\times 10^{7}.$ This is a surprisingly large number. Moreover, $0.91-\epsilon(x_{0})\approx 0.997951789>\beta=$ 0.9 . The reason for such results is that $T$ is not a contraction mapping. Alternatively, one can find $x_{1}\in(0,1)$ that minimises $1-\epsilon(x)$ over $(0,1)$ and then calculate $B(x_{1}).$ Then
$x_{1}=\frac{79-8\sqrt{94}}{3}\approx 0.47904076,\quad\mbox{and}$
$B(x_{1})=\frac{3(377+4\sqrt{94})}{1250}\approx 0.997875453.$
Note that $1-\epsilon(x_{1})<1-\epsilon(x_{0}),$ but $B(x_{1})(1-\epsilon(x_{1}))>B(x_{0})(1-\epsilon(x_{0})).$ The operator $F_{3}$ proposed in Subsection 2.3 to study value iterations is
$F_{3}v(s)=\min_{a\in A(s)}\left(r(s,a)+\beta\sqrt{\int_{S}v^{2}(s^{\prime})q(ds^{\prime}|s,a)}\right)$
and is a $\beta$-contraction. Since the algorithm using the operator $F_{3}$ can start with $w_{0}=\hat{0},$ the constant $L$ in (34) is equal to $505/(1-\beta)=$ 50 . The difference between the algorithm using Du’s theorem for concave and increasing operators and our proposal is significant. One should remember that the solution to Bellman equation (4) in this example is the limit $\lim_{n\to\infty}w_{n}^{1/(1-\rho)}=w_{*}^{-4}.$
Remark 20
We would like to point out that in our method of studying existence and uniqueness of a solution to the Bellman equation we do not need boundary conditions mentioned in Remark 19. Therefore, we can drop the assumption made in the literature in case when $\gamma,\ \rho>1$ that the per-period utilities $u(s,a)\geq u_{0}$ for some $u_{0}>0$ and for all $(s,a)\in\mathbb{K}.$ However, if $\inf_{s\in S}\min_{a\in A(s)}u(s,a)=0,$ then $r(s,a)=(1-\beta)u(s,a)^{1-\rho}$ is unbounded and there should be a weight function $\omega$ satisfying extra assumptions.
5 Conclusions
We consider dynamic programming problems arising from Markov decision processes with Epstein-Zin preferences involving certainty equivalents induced by power functions. The values of parameters $\gamma$ and $\rho$ are such that the aggregators that we deal with are according to Marinacci and Montrucchio (2010) of Thompson type. We have shown that in the two cases: (1) $10<\rho<\gamma<$ 1 and (2) $1<\rho<\gamma$ studied in Ren and Stachurski (2020, 2021) one can use the Banach contraction mapping principle instead of a fixed point theorem of Du (1990) for increasing and convex or concave operators acting on an ordered Banach space. The contraction mapping theorem is applied to some sets of non-negative functions defined on the state space, endowed with the standard weighted supremum metric. This is a much simpler way for studying the existence and uniqueness of recursive utilities for the models with Epstein-Zin aggregators than the method using the Thompson metrics as in Marinacci and Montrucchio (2010) or the fixed point theorem of Du (1990) as in Ren and Stachurski (2020, 2021). Our assumptions on the per-period utility functions are weaker than the corresponding ones in Marinacci and Montrucchio (2010) and Ren and Stachurski (2020, 2021). For example, we do not assume the boundary conditions assuring that the fixed point of the operator involved in the analysis lies in the interior of its domain. In dynamic programming the boundary conditions are connected with assumptions (25) and (26) in Proposition 18. The constant $\epsilon>0$ in (25) and (26) has an influence on the speed of convergence of the dynamic programming operator to a fixed point. Using the Banach theorem, we do not need to apply any boundary conditions and therefore, we are able to provide better estimation for the convergence of iterations of operators used in modified value iteration process. The relevance of this result is that our method requires a smaller number of iterations needed to reach a given accuracy of the solution to the Bellman equation than in the case of applying Du’s theorem. It seems that the third case considered by Ren and Stachurski (2020) where $0<\rho<1<\gamma$ cannot be analysed by Banach’s theorem. However, we wish to point out that our techniques apply to other two cases (1) $0<\gamma<\rho$ and (2) $1<\gamma<\rho,$ where the theorem of Du (1990) is inapplicable since the dynamic programming operators are then neither convex nor concave.