### v2

# Stochastic dynamic programming under recursive Epstein-Zin preferences

Abstract This paper investigates discrete-time Markov decision processes with recursive utilities (or payoffs) defined by the classic CES aggregator and the Kreps-Porteus certainty equivalent operator. According to the classification introduced by Marinacci and Montrucchio, the aggregators that we consider are Thompson. We focus on the existence and uniqueness of a solution to the Bellman equation. Since the per-period utilities can be unbounded, we work with the weighted supremum norm. Our paper shows three major points for such models. Firstly, we prove that the Bellman equation can be obtained by the Banach fixed point theorem for contraction mappings acting on a standard complete metric space. Secondly, we need not assume any boundary conditions, which are present when the Thompson metric or the Du’s theorem are used. Thirdly, our results give better bounds for the geometric convergence of the value iteration algorithm than those obtained by Du’s fixed point theorem. Moreover, our techniques allow to derive the Bellman equation for some values of parameters in the CES aggregator and the Kreps-Porteus certainty equivalent that cannot be solved by Du’s theorem for increasing and convex or concave operators acting on an ordered Banach space.
JEL classification: C61, C62, D81, E20
Keywords: Recursive utility; Thompson aggregator; Bellman equation; Dynamic programming
1 Introduction
In the classical setting the additive utility is constructed with the aid of a linear aggregator of a current utility (or reward) and the expected discounted utility from tomorrow onwards. More precisely, if $c=(c_{n})$ is a consumption sequence and $u$ is the per-period utility, then the utility $\widehat{V}_{n}(c)$ from period $n$ onwards solves the forward-looking equation
$\widehat{V}_{n}(c)=u(c_{n})+\beta\mathbb{E}_{n}(\widehat{V}_{n+1}(c))\quad\mbox{and}\quad\widehat{V}_{n}(c)=\mathbb{E}_{n}\left(\sum_{k\geq n}\beta^{k-n}u(c_{k})\right).$
Here, $\beta\in[0,1)$ is a discount factor and $\mathbb{E}_{n}$ is the expectation operator with respect to the time period $n$. Although this framework is mathematically attractable, it does not take into account the risk associated with transitions from state to state. In order to include the risk aversion into a multi-period decision process, we replace the expected discounted utility by a certainty equivalent operator induced by an increasing concave utility function defined on the set of non-negative real numbers. One of the possibilities is to choose a certainty equivalent of the exponential function. Its negative is sometimes called an entropic risk measure and gains much popularity in economics, see Miao (2014); Sargent and Stachurski (2024). Hansen and Sargent (1995) were among first who applied such a framework to a linear quadratic Gaussian control problem. Their approach inspired research in other areas, e.g., in stochastic optimal growth theory (see Bäuerle and Jaśkiewicz (2018)) and in Markov controlled models (see Asienkiewicz and Jaśkiewicz (2017)).
Following the seminal work of Kreps and Porteus (1978), Epstein and Zin (1989) suggested a recursive utility framework that permits for a separation between risk attitudes and the degree of intertemporal substitution. It generalises a deterministic recursive utility introduced by Koopmans (1960) to a stochastic setting. This framework can also allow for early or late resolution of uncertainty and is tractable to apply because the dynamic programming technique can be used. More precisely, Epstein and Zin (1989) define a recursive utility under uncertainty by two primitives: a time aggregator and a conditional certainty equivalent. Then, the recursive utility $V_{n}(c)$ from the period $n$ onwards can be written as follows
$V_{n}(c)=\left[(1-\beta)u(c_{n})^{1-\rho}+\beta{\cal M}_{n}(V_{n+1}(c))^{1-\rho}\right]^{\frac{1}{1-\rho}},$
where $10<\rho\not=$ 1 . Here, ${\cal M}_{n}$ is a certainty equivalent and can be induced for instance by the power function:
${\cal M}_{n}(V_{n+1}(c))=\mathbb{E}_{n}(V_{n+1}(c)^{1-\gamma})^{\frac{1}{1-\gamma}},\quad 0<\gamma\not=1.$
This certainty equivalent is also known under the name of the Kreps-Porteus operator, see Miao (2014). The value $1/\rho$ represents the elasticity of intertemporal substitution between the composite good and the certainty equivalent, whereas $\gamma$ governs the level of relative risk aversion with respect to atemporal gambles. Therefore, this time aggregator is known under the name of CES (Constant Elasticity of Substitution) aggregator. Recursive preferences of this type are called the Epstein-Zin preferences. They form a core of the quantitative asset pricing literature and find applications ranging from optimal taxation to fiscal policy and business cycles, see Bansal and Yaron (2004), Kaplan and Violante (2014) and Schorfheide et al. (2018).
The growing popularity of recursive preferences gives rise to the expanding domain of their applications in macroeconomics and finance. The reader is referred in this matter, among others, to Backus et al. (2005); Miao (2014); Sargent and Stachurski (2024); Skiadas (2009); xing. Moreover, the predominant acceptance of recursive preferences has recently led to a stream of fundamental issues such as existence, convergence and uniqueness. For instance, Hansen and Scheinkman (2012) established a connection between the recursive utility equation and the spectral radius of a corresponding valuation operator. Their approach fostered to further developments, where Borovićka and Stachurski (2020) showed that the spectral radius condition is necessary and sufficient for existence of the Epstein-Zin recursive utility. Finally, Christensen (2022) making use of this method provided a characterisation of recursive utility under risk sensitivity, ambiguity and Epstein-Zin preferences. Uniqueness can be obtained only under restrictions on the state space and the Markov transition. However, none of the aforementioned papers studies dynamic programming with recursive utility.
Despite the substantial progress on recursive utility, less is known about its implications for dynamic programming. Dynamic programming with recursive utility in economics was commenced by Streufert (1990) and Ozaki and Streufert (1996), who introduced a notion of biconvergence. This condition means that utility values can be approximated by increasing and decreasing orbits. Recently, along similar lines, Bich et al. (2018) studied deterministic recursive programs under some regularity properties. It is worthy to mention that the work of Denardo (1967) was one of the first papers that are related to dynamic programming with recursive preferences.
Marinacci and Montrucchio (2010, 2019) introduce a contraction-type approach to recursive utilities. More precisely, Marinacci and Montrucchio (2010) divided the Epstein-Zin aggregators into two classes. In one of them the aggregators are called Blackwell and the existence and uniqueness of recursive utility can be proved by the Banach contraction mapping principle. This approach corresponds to the work of Blackwell (1965) on discounted dynamic programming. In the second class, the aggregators are called Thompson, for which the standard methods and metrics cannot be directly applied. The concave Epstein-Zin aggregators are also of Thompson-type. Marinacci and Montrucchio (2010) exploited recursive utilities derived from the second class and obtained significant results by applying the Thompson fixed point theorem for non-linear operators on ordered normed spaces with a complete and normal positive cone (consult with Thompson (1963)). Another approach to Thompson aggragators has been recently proposed by Becker and Rincón-Zapatero (2023), where uniqueness of recursive utility is presented differently than in Marinacci and Montrucchio (2010, 2019).
Dynamic programming with monotone concave and $\alpha$-concave operators was recently adapted by Bloise and Vailakis (2018) and Balbus (2020), respectively. They demonstrate new results on the existence and uniqueness of a recursive utility as well as some applications to Bellman equations in dynamic optimisation. The concave and convex dynamic programming operators have been further developed by Ren and Stachurski (2020, 2021). They obtained valuable results on the existence and uniqueness of solutions of the Bellman equation using a fixed point theorem of Du (1990) for monotone convex or concave operators. They examined the Epstein-Zin preferences with unbounded rewards. However, this promising approach based either on the Thompson metric or Du’s theorem is frustrated by the fact that uniqueness only obtains in the interior of the domain, or subject to some appropriate boundary condition. In many relevant applications, this sort of boundary condition is unnatural. For instance, it imposes restrictions on per-period utilities, and moreover, it has an influence on the speed of convergence of iterations of dynamic programming operators to the value function. A unified treatment of recursive methods under minimal assumptions on the Koopmans aggregator has been recently proposed by Bloise et al. (2024). In particular, they developed an elementary approach to dynamic programming with recursive utility based on Tarski’s fixed point theorem and establish the existence of a value of the recursive program.
The structure of non-linear recursive preferences turned out to be harder to study. It was realised and emphasised in several papers that the operators involved in the construction of recursive utility are not contractive in the supremum norm. Consequently, the Epstein-Zin preferences fail to satisfy the traditional Blackwell’s discounting condition and an appeal to the Banach contraction mapping principle is of limited scope. Even establishing the existence of a recursive utility function has seemed to be an arduous task. In this paper, we show that it is not true. Our results for Epstein-Zin preferences are presented within an $MDP$ (Markov decision processes) framework. It worth emphasising that $MDP$ models play an important role in operations research, economics and finance, see Acemoglu (2009); Bäuerle and Rieder (2011); Hernández-Lerma and Lasserre (1996, 1999); Kochenderfer (2015); Miao (2014); Stachurski (2009).
Our work is inspired by the results of Ren and Stachurski (2020, 2021), who applied Du’s theorem to solve an $MDP$ model with the Epstein-Zin preferences in three cases: $0<\rho<\gamma<1,$ $1<\rho<\gamma$ and $0<\rho<1<\gamma.$ As noted by Bansal and Yaron (2004); Schorfheide et al. (2018) the case $\rho<\gamma$ is most empirically relevant implying that the agent prefers early resolution of uncertainty. We allow the per-period utilities to be unbounded and use the weighted supremum norms. The recursive utility is then a fixed point in a suitably chosen function space. We prove the existence and uniqueness of a solution to the Bellman equation in the first two aforementioned cases using the Banach contraction mapping principle for the space of upper semicontinuous functions endowed with the weighted supremum norm, see Wessels (1977). This method works for some classes with convex aggregators that are according to Marinacci and Montrucchio (2010) of Thompson type. Using the Banach fixed point theorem and the supremum norm we get a simple estimation of the speed of geometric convergence of the value iteration algorithm to a solution to the Bellman equation. As a starting point we can choose a constant function, in one case the zero function. We give two examples with bounded rewards showing that our value iterations converge faster than in Ren and Stachurski (2020, 2021) (that is based on Du’s theorem). Moreover, our regularity assumptions on the per-period utility function and inequalities involving the weight function are weaker than those used in Ren and Stachurski (2020, 2021). We also show the existence and uniqueness of the solution to the Bellman equation, with some values of parameters in the Epstein-Zin preferences, where the dynamic programming operators are neither convex nor concave. The fixed point theorem of Du (1990) in these cases is useless. An application of the Banach contraction principle to dynamic programming in $MDPs$ with the Epstein-Zin preferences is possible due to some simple transformations of the Bellman equations.
The paper is organised as follows. In Section 2 we provide preliminaries and present our major theorems. We start with models, where the agent is allowed to use only stationary policies. All proofs from Subsection 2.3 are collected in Section 3. Next in Subsection 2.4 we extend our analysis to a non-stationary setting, where history-dependent policies are permitted. In Section 4 we compare our assumptions imposed on the primitives with the ones accepted in Ren and Stachurski (2020, 2021). We also provide examples showing that the Banach contraction mapping principle gives better bounds for convergence of value iteration algorithm than Du’s theorem. Finally, Section 5 summarises our main points in this paper.
2 Dynamic programming model and main results
2.1 Preliminaries
[…]
$\|v\|_{\omega}:=\sup_{s\in S}\frac{|v(s)|}{\omega(s)}.$
By ${\cal C}^{\omega}$ (${\cal U}^{\omega}$) we denote the set of all continuous (upper semicontinuous) functions on $S$ for which $\|\cdot\|_{\omega}$-norm is finite and by ${\cal C}_{+}^{\omega}$ (${\cal U}_{+}^{\omega}$) we denote the set of non-negative functions in ${\cal C}^{\omega}$ (${\cal U}^{\omega}$).
Assume that $(X,d)$ is a complete metric space and $\delta\in(0,1).$ A mapping $F:X\to X$ is called a $\delta$-contraction, if $d(Fx_{1},Fx_{2})\leq\delta d(x_{1},x_{2})$ for all $x_{1},x_{2}\in X.$ By the Banach contraction mapping principle $F$ has a unique fixed point $x_{*}\in X.$ Moreover, for any element $x_{0}\in X$
$d(F^{n}x_{0},x_{*})\leq\frac{\delta^{n}}{1-\delta}d(Fx_{0},x_{0}),$ (1)
where $F^{n}$ is the $n$-th iteration of $F$, see Section 3.12 in Aliprantis and Border (2006).
Let $Z$ a separable metric space. By $\varphi$ we denote a correspondence from $S$ to $Z.$ For any $D\subset Z$ we put
$\varphi^{-1}(D):=\{s\in S:\varphi(s)\cap D\not=\emptyset\}.$
If $\varphi^{-1}(D)$ is a closed (an open) subset in $S$ for each closed (open) subset $D$ of $Z,$ then $\varphi$ is said to be upper (lower) semicontinuous. If $\varphi$ is both upper and lower semicotinuous, then it is called continuous.
2.2 A model with stationary recursive utilities
In this paper we study a Markov decision process defined by the following objects.
  • (i)
    $S$ is a set of states and is assumed to be a Borel space.
  • (ii)
    $A$ is the action space and is also assumed to be a Borel space.
  • (iii)
    $\mathbb{K}$ is a non-empty Borel subset of $S\!\times\!A$. We assume that for each $s\in S$, the non-empty $s$-section
    $A(s)=\{a\in A:(s,a)\in\mathbb{K}\}$
    of $\mathbb{K}$ represents the set of available actions in state $s$.
  • (iv)
    $q$ is a Feller transition probability from $\mathbb{K}$ to $S,$ i.e., the function
    $(s,a)\to\int_{S}v(s^{\prime})q(ds^{\prime}|s,a)$
    is continuous for any $v\in{\cal C}.$
  • (v)
    $u:\mathbb{K}\to\mathbb{R}$ is a upper semicontinuous non-negative per-period utility function.
  • (vi)
    $\beta\in[0,1)$ is a discount factor.
We first study a stationary model in which an agent uses the same decision rule in every period. Let $\Phi$ be the set of all Borel measurable mappings $f$ from $S$ into $A$ such that $f(s)\in A(s)$ for each $s\in S$. By Theorem 1 in Brown and Purves (1973), $\Phi\not=\emptyset.$ A stationary policy is a constant sequence $\pi=(f,f,\ldots)$ where $f\in\Phi$ and therefore can be identified with the Borel mapping $f.$
Assume for a moment that $u$ is bounded. A state-action aggregator $H$ maps feasible state-action pairs $(s,a)\in\mathbb{K}$ and bounded Borel measurable functions $w$ into real values $H(s,a,w)$ representing lifetime utility, contingent on the current action $a$, the current state $s$ and the use of $w$ to evaluate future states. Traditional additively separable case is implemented by setting:
$H(s,a,w)=u(s,a)+\beta\int_{S}w(s^{\prime})q(ds^{\prime}|s,a),$
where $(s,a)\in\mathbb{K}.$ Under Epstein-Zin preferences, the state-action aggregator $H$ is of the form
$H(s,a,w)=\left[r(s,a)+\beta\left(\int_{S}w(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},\ \ (s,a)\in\mathbb{K},$ (2)
[…]
$r(s,a)=(1-\beta)u(s,a)^{1-\rho},\quad(s,a)\in\mathbb{K}.$ (3)
Definition 1
A function $\widetilde{v}$ is a solution to the Bellman equation, if
$\widetilde{v}(s)=\max_{a\in A(s)}H(s,a,\widetilde{v})\quad\mbox{for all}\quad s\in S.$ (4)
Definition 2
For any policy $f\in\Phi$, we call $v_{f}$ an $f$-value function, if it satisfies the equation
$v_{f}(s)=H(s,f(s),v_{f}),\quad s\in S$
with $H$ defined in (2).
The value function $v_{f}$ is a recursive utility in the infinite horizon model in the sense of Koopmans (1960), see also Marinacci and Montrucchio (2010). It is determined by the aggregator (2) and a policy $f\in\Phi.$
Definition 3
A stationary policy $f_{*}\in\Phi$ is optimal, if
$v_{f_{*}}(s)=\widetilde{v}(s)\geq v_{f}(s)\quad\mbox{for all }f\in\Phi\mbox{ and }s\in S,$
and $\widetilde{v}$ is a solution to the Bellman equation (4).
Such a restriction to the class of stationary policies and the recursive utilities in the sense of Koopmans (1960) is also used by Ren and Stachurski (2020, 2021).
If there exists a policy $f\in\Phi$ that maximises the right-hand side of (4), then, as we show in the sequel, $f$ turns out to be an optimal stationary policy. This case is well-known in the literature with standard affine aggregators and expected value as certainty equivalent, see Bertsekas (2022); Hernández-Lerma and Lasserre (1996).
2.3 Main results on optimal policies in stationary models
The Bellman equation in (4) is precisely of the form
$\widetilde{v}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}\widetilde{v}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},$ (5)
for all $s\in S,$ provided that the integral of $\widetilde{v}(s)^{1-\gamma}$ exists with respect to $q(\cdot|s,a)$ for every $(s,a)\in\mathbb{K}.$ Define
$\theta:=\frac{1-\gamma}{1-\rho}.$
This number will be used in all four cases of parametrisation in the Epstein-Zin preferences considered in this paper. Let $\hat{0}$ be the zero function.
In order to find a solution $v$ to the Bellman equation (5) when $\gamma,\ \rho\in(0,1)$, we first formulate the following conditions.
  • (A1)
    For each $s\in S$, the set $A(s)$ is compact and, moreover, the correspondence $s\to A(s)$ is upper semicontinuous.
  • (A2)
    There exists a continuous function $\omega:S\to[1,\infty)$ such that
    $r(s,a)\leq M\omega(s),\quad(s,a)\in\mathbb{K}$
    for some constant $M>0.$
  • (A3)
    The function
    $(s,a)\to\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)$
    is continuous and there is a constant $c\geq 1$ such that
    $\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)\leq c\omega(s)$
    for all $(s,a)\in{\mathbb{K}}$ and $c\beta^{\theta}<1.$
  • (A4)
    The function
    $(s,a)\to\int_{S}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)$
    is continuous on $\mathbb{K}$ and there is a constant $c\geq 1$ such that
    $\int_{S}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\leq c\omega(s)^{\theta}$
    for every $(s,a)\in{\mathbb{K}}$ and $c^{\frac{1}{\theta}}\beta<1.$
Case 1: $10<\rho<\gamma<$ 1 .
Note that $10<\theta<$ 1 and for any function $V\in{\cal U}_{+}^{\omega}$, define the operator $F_{1}$ as follows:
$F_{1}V(s):=\max_{a\in A(s)}H_{1}(s,a,V),\ \ s\in S,$
where
$H_{1}(s,a,V):=\left[r(s,a)+\beta\left(\int_{S}V(s^{\prime})q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]^{\theta},\ \ (s,a)\in\mathbb{K}.$ (6)
We can now state our first main result.
Theorem 1
Consider Case 1. Under assumptions (A1)-(A3), the following assertions hold.
$(a)$ The Bellman equation (4) has a unique solution $v_{*}\in{\cal U}_{+}^{\omega}$.
$(b)$ The mapping $F_{1}:{\cal U}_{+}^{\omega}\to{\cal U}_{+}^{\omega}$ is a $c\beta^{\theta}$-contraction. It has a unique fixed point $w_{*}$ and $v_{*}=w_{*}^{\frac{1}{1-\gamma}}$. Moreover, if $F^{n}_{1}$ is the $n$-th composition of $F_{1}$ with itself, then
$\|F_{1}^{n}\hat{0}-w_{*}\|_{\omega}\leq\frac{c^{n}\beta^{n\theta}}{1-c\beta^{\theta}}M.$ (7)
$(c)$ There exists $f_{*}\in\Phi$ that satisfies the optimality equation
$v_{*}(s)=H(s,f_{*}(s),v_{*})\quad\mbox{for all}\quad s\in S.$
$(d)$ The stationary policy $f_{*}$ is optimal.
Case 2: $10<\gamma<\rho<$ 1 .
Note that $\theta>1$ and for any function $V\in{\cal U}_{+}^{\omega}$, define the operator $F_{2}$ as follows:
$F_{2}V(s):=\max_{a\in A(s)}H_{2}(s,a,V),\ \ s\in S,$
where
$H_{2}(s,a,V):=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}V(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right],\ \ (s,a)\in\mathbb{K}.$
We assume the integrability of the function $V(\cdot)^{\theta}$ with respect to $q(\cdot|s,a)$ for $(s,a)\in\mathbb{K}.$
We can now state our second main result.
Theorem 2
Consider Case 2. Under assumptions (A1)-(A2) and (A4), the following assertions hold.
$(a)$ The Bellman equation (4) has a unique solution $v_{*}\in{\cal U}_{+}^{\omega}$.
$(b)$ The mapping $F_{2}:{\cal U}_{+}^{\omega}\to{\cal U}_{+}^{\omega}$ is a $c^{\frac{1}{\theta}}\beta$-contraction. It has a unique fixed point $w_{*}$ and $v_{*}=w_{*}^{\frac{1}{1-\rho}}$. Moreover, if $F^{n}_{2}$ is the $n$-th composition of $F_{2}$ with itself, then
$\|F_{2}^{n}\hat{0}-w_{*}\|_{\omega}\leq\frac{c^{\frac{n}{\theta}}\beta^{n}}{1-c^{\frac{1}{\theta}}\beta}M.$ (8)
$(c)$ There exists $f_{*}\in\Phi$ that satisfies the optimality equation
$v_{*}(s)=H(s,f_{*}(s),v_{*})\quad\mbox{for all}\quad s\in S.$
$(d)$ The stationary policy $f_{*}$ is optimal.
If $\gamma,\ \rho>1,$ then we must make some changes in our assumptions.
  • (B1)
    For each $s\in S$, the set $A(s)$ is compact and, moreover, the correspondence $s\to A(s)$ is continuous.
  • (B2)
    There exists a continuous function $\omega:S\to[1,\infty)$ such that
    $r(s,a)\leq M\omega(s),\ \ \mbox{for all}\ \ (s,a)\in\mathbb{K}$
    and for some constant $M>0.$ The function $r$ is positive and continuous on $\mathbb{K}.$
  • (B3)
    The function
    $(s,a)\to\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)$
    is continuous and there is a constant $c\geq 1$ such that
    $\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)\leq c\omega(s)$
    for all $(s,a)\in{\mathbb{K}}$ and $c\beta^{\theta}<1.$
  • (B4)
    The function
    $(s,a)\to\int_{S}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)$
    is continuous on $\mathbb{K}$ and there is a constant $c\geq 1$ such that
    $\int_{S}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\leq c\omega(s)^{\theta}$
    for every $(s,a)\in{\mathbb{K}}$ and $c^{\frac{1}{\theta}}\beta<1.$
The difference between (A3) and (B3) and between (A4) and (B4) lies in the properties of the set $\mathbb{K}.$
Case 3: $\gamma>\rho>1.$
Now we have $\theta>1.$ Therefore, the Bellman equation in (4) has the following equivalent form:
$w(s)=\min_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}w(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]$
for all $s\in S.$ For any function $V\in{\cal C}_{+}^{\omega}$, define the operator $F_{3}$ as follows:
$F_{3}V(s):=\min_{a\in A(s)}H_{3}(s,a,V),\ \ s\in S,$
where
$H_{3}(s,a,V):=r(s,a)+\beta\left(\int_{S}V(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}$
for all $(s,a)\in\mathbb{K}.$ We shall show that $F_{3}$ maps ${\cal C}_{+}^{\omega}$ into itself. Our next result is as follows.
Theorem 3
Consider Case 3. Under assumptions (B1)-(B2) and (B4), the following assertions hold.
$(a)$ The Bellman equation (4) has a unique positive solution $v_{*}\in{\cal C}_{+}^{\omega}$.
$(b)$ The mapping $F_{3}:{\cal C}_{+}^{\omega}\to{\cal C}_{+}^{\omega}$ is a $c^{\frac{1}{\theta}}\beta$-contraction. It has a unique fixed point $w_{*}$ and $v_{*}=w_{*}^{\frac{1}{1-\rho}}$. Moreover, if $F^{n}_{3}$ is the $n$-th composition of $F_{3}$ with itself, then
$\|F_{2}^{n}\hat{0}-w_{*}\|_{\omega}\leq\frac{c^{\frac{n}{\theta}}\beta^{n}}{1-c\beta}M.$ (9)
$(c)$ There exists $f_{*}\in\Phi$ that satisfies the optimality equation
$v_{*}(s)=H(s,f_{*}(s),v_{*})\quad\mbox{for all}\quad s\in S.$
$(d)$ The stationary policy $f_{*}$ is optimal.
Case 4: $\rho>\gamma>1.$
Now we have $\theta\in(0,1),$ but $01-\gamma<$ 0 and $01-\rho<$ 0 . Define the operator $F_{4}:{\cal C}_{+}^{\omega}\to{\cal C}_{+}^{\omega}$ as follows:
$F_{4}V(s):=\min_{a\in A(s)}H_{4}(s,a,V),\ \ s\in S,$
where
$H_{4}(s,a,V):=\left[r(s,a)+\beta\left(\int_{S}V(s^{\prime})q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]^{\theta},\ \ (s,a)\in\mathbb{K}.$
Now we state our last main result in this section.
Theorem 4
Consider Case 4. Under assumptions (B1)-(B3), the following assertions hold.
$(a)$ The Bellman equation (4) has a unique solution $v_{*}\in{\cal C}_{+}^{\omega}.$
$(b)$ The mapping $F_{4}:{\cal C}_{+}^{\omega}\to{\cal C}_{+}^{\omega}$ is a $c\beta^{\theta}$-contraction. It has a unique fixed point $w_{*}$ and $v_{*}=w_{*}^{\frac{1}{1-\gamma}}$. Moreover, if $F^{n}_{4}$ is the $n$-th composition of $F_{4}$ with itself, then
$\|F_{4}^{n}\hat{0}-w_{*}\|_{\omega}\leq\frac{c^{n}\beta^{n\theta}}{1-c\beta^{\theta}}M.$ (10)
$(c)$ There exists $f_{*}\in\Phi$ that satisfies the optimality equation
$v_{*}(s)=H(s,f_{*}(s),v_{*})\quad\mbox{for all}\quad s\in S.$
$(d)$ The stationary policy $f_{*}$ is optimal.
The proofs of Theorems 1-4 are given in Section 3.
Remark 1
$(a)$ If the function $r$ is bounded, then in our assumptions we can take $\omega(x)\equiv 1$ and $c=1.$
$(b)$ The class of functions $r$ and transition probabilities satisfying our assumptions involving the weight function $\omega$ is pretty large. For various examples the reader is referred to Durán (2000) and Jaśkiewicz and Nowak (2011). Here, we would like to show to construct of an appropriate function $\omega.$ It is quite easy to find a continuous function $\widehat{\omega}:S\to[1,\infty)$ such that
$r(s,a)\leq M\widehat{\omega}(s)\ \ \mbox{for all}\ \ s\in S.$
For any $\beta\in(0,1)$ one can find $\alpha>1$ such that $\alpha\beta<1.$ Suppose that the difference
$D(s,a):=\int_{S}\widehat{\omega}(s^{\prime})q(ds^{\prime}|s,a)-\alpha\widehat{\omega}(s)>0$
on some subset $\mathbb{K}_{0}$ of $\mathbb{K}$ and suppose that $D(s,a)$ is bounded from above. Let $z>0$ and $\omega(s):=z+\widehat{\omega}(s),$ $s\in S.$ Then
$\int_{S}\omega(s^{\prime})q(ds^{\prime}|s,a)=\alpha\omega(s)+(1-\alpha)z+D(s,a).$
If we take $z$ such that $(1-\alpha)z+D(s,a)<0$ for all $(s,a)\in\mathbb{K}_{0},$ then
$\int_{S}\omega)s^{\prime})q(ds^{\prime}|s,a)\leq\alpha\omega(s)\ \ \mbox{for all}\ \ (s,a)\in\mathbb{K}.$
In this way, we can establish an appropriate weight function for several cases arising in economic dynamics.
Remark 2
From (B2) it follows that
$r_{0}(s):=\min_{a\in A(s)}r(s,a)>0\ \ \mbox{ for all}\ \ s\in S.$
Hence, the operators $F_{3}$ and $F_{4}$ are considered on ${\cal C}_{+}^{\omega}.$ In particular, $F_{3}\hat{0}$ and $F_{4}\hat{0}$ are well-defined. Therefore, the fixed point of $F_{i}$ (for $i=3,4$) can be obtained by iterating the operator on the zero function, which makes the approximation procedure simpler.
Remark 3
Cases 1 and 3 were considered in Ren and Stachurski (2020, 2021). Their proof of the existence of a unique solution to the Bellman equation is based on the fixed point theorem of Du (1990) for monotone convex or concave operators on ordered Banach spaces. They also show the geometric convergence of iterations of the corresponding operators. However, this speed of convergence is slower than in Theorems 1 and 3. We discuss this issue in Subsection 4.2. The results of Ren and Stachurski (2020, 2021) also apply to the case when $0<\gamma<1<\rho$ which is not covered by our approach. On the other hand, Cases 2 and 4 cannot be solved by using the theorem of Du (1990), since the corresponding operators are in these aforementioned cases neither concave nor convex. For more comments on this issue see Subsection 4.1.
Remark 4
Note that our conditions in Theorem 3 are weaker than those in Ren and Stachurski (2020), who assumed that the reward function is bounded from below by some function greater than 1. Moreover, they also impose that the integral of the weight function is bounded from below by this weight function multiplied by some constant smaller than $1/\beta^{\theta}.$ The reader is referred to Assumption 4.1 and inequalities (42) and (44) in Ren and Stachurski (2020).
Remark 5
In the most empirically relevant parametric specifications $\rho<\gamma$, implying that the agent prefers early resolution of uncertainty (Cases 1 and 3). For a discussion of this topic the reader is referred to Bansal and Yaron (2004) or Schorfheide et al. (2018). Cases 2 and 4 where the coefficient of relative risk aversion $\gamma>1$ and the intertemporal elasticity of substitution $\frac{1}{\rho}<1$ are also relevant, see Bansal and Yaron (2004); Basu and Bundick (2017); Farhi and Werning (2008).
2.4 An extension to models with non-stationary policies
A definition of the decision process involving the Epstein-Zin aggregators in a framework of general class of history dependent policies is not obvious. For some specifications of the parameters $\gamma$ and $\rho$, one can define utilities in finite step models and then consider the limit as the time horizon tends to infinity. Then using iterations of some appropriately defined dynamic programming operators, one can prove that an optimal stationary policy derived from the Bellman equation is also optimal in the class of all policies.
A similar construction in risk-sensitive Markov decision processes involving the additive (affine) aggregator and entropic risk measure as a certainty equivalent is described in Asienkiewicz and Jaśkiewicz (2017); Bäuerle and Jaśkiewicz (2018). Additional comments on this issue can be found in Sargent and Stachurski (2024).
A policy is a sequence of Borel measurable mappings $\pi=(\pi_{n})_{n\in\mathbb{N}}$ from the history space to the action set. More precisely, each $\pi_{n}(h_{n})\in A(s_{n})$, $n\in\mathbb{N},$ where $h_{n}=(s_{1},a_{1},\ldots,s_{n-1},a_{n-1},s_{n})$ is the history of the process up to the $n$-th state. For $n=1,$ $h_{1}=s_{1}.$ The class of all policies is denoted by $\Pi$. A Markov policy is a sequence $\pi=(\pi_{n})_{n\in\mathbb{N}}$ where each $\pi_{n}$ depends only on the $n$-th state $s_{n}\in S.$ The class of all Markov policies is denoted by $\Pi_{M}$.
To avoid complex notation, we shall confine ourselves to consideration of Markov policies. The proof that Markov policies are enough in dynamic programming is standard and will be omitted in this paper. With respect to this issue consult with Bertsekas (2022); Hernández-Lerma and Lasserre (1996); Sargent and Stachurski (2024).
Let us fix a Markov policy $\pi=(\pi_{n})_{n\in\mathbb{N}}$. For a non-negative function $\psi:S\to\mathbb{R}$ define the operator
${\cal T}_{\pi_{n}}\psi(s_{n}):=\left[r(s_{n},\pi_{n}(s_{n}))+\beta\left(\int_{S}\psi(s_{n+1})^{1-\gamma}q(ds_{n+1}|s_{n},\pi_{n}(s_{n}))\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}}$ (11)
for all $s_{n}\in S.$ We assume again the integrability of the function $\psi(\cdot)^{1-\gamma}$ with respect to $q(\cdot|s,a)$ for every $(s,a)\in\mathbb{K}.$ If $1-\gamma<0,$ then $\psi$ must be positive. Recall that $r$ is defined in (3). The operator in (11) is monotone for some values of parameters $\gamma$ and $\rho$. Here, we shall consider cases discussed in the previous section.
Case (A): $\gamma,\ \rho\in(0,1).$ Moreover, (A1)-(A2) are satisfied and if $\theta\in(0,1)$, then (A3) holds and if $\theta>1,$ then (A4) holds.
The recursive utility in the finite time horizon is defined by the following composition of the operators
$U_{n}(\pi)(s_{1}):={\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\hat{0}(s_{1}).$ (12)
Here, the process starts at state $s_{1}\in S$ and terminates after $n$ steps. with the terminal utility which equals zero. Since $\gamma,\ \rho\in(0,1),$ it follows that $U_{m}(\pi)(s_{1})\leq U_{m+1}(\pi)(s_{1})$ for all $m\geq 1.$ The recursive utility of an agent in the infinite time horizon is defined as
$U(\pi)(s_{1}):=\lim_{m\to\infty}U_{m}(\pi)(s_{1}),\ \ s_{1}\in S,\ \pi\in\Pi_{M}.$
Definition 4
A policy $\pi_{*}$ is optimal, if
$U(\pi_{*})(s_{1})\geq U(\pi)(s_{1})\quad\mbox{for all}\quad\pi\in\Pi,\ s_{1}\in S.$
The composition ${\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\psi(s_{1})$ can be defined for any $\psi\geq 0$ and we have
${\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\psi(s_{1})\geq{\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\hat{0}(s_{1}).$
[…]
$U(f)(s_{1})=H(s_{1},f(s_{1}),U(f))\ \ \mbox{for all }\ \ s_{1}\in S,$ (13)
where $H$ is given in (2). Let us define
${\cal T}\psi(s_{n}):=\max_{a_{n}\in A(s_{n})}H(s_{n},a_{n},\psi),\ \ \psi\geq 0.$ (14)
Note that the sequence ${\cal T}^{m}\hat{0}$ of $m$ compositions of the operator ${\cal T}$ with itself is non-decreasing and for any Markov policy $\pi$, we have
$U(\pi)(s_{1})=\lim_{m\to\infty}{\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{m}}\hat{0}(s_{1})\leq\lim_{m\to\infty}{\cal T}^{m}\hat{0}(s_{1})\ \ \mbox{for all}\ \ s_{1}\in S.$
Theorem 5
In Case (A), there exists a stationary optimal policy in the class of all policies.
Proof By Theorem 1 or 2 (depending on whether (A3) or (A4) is assumed), the Bellman equation has a unique solution $v_{*}$ (in appropriate function space) and there exists $f_{*}\in\Phi$ such that
$v_{*}(s_{1})={\cal T}v_{*}(s_{1})=H(s_{1},f_{*}(s_{1}),v_{*})\quad\mbox{for all}\quad s_{1}\in S.$
By (13) we also have $U(f_{*})(s_{1})=H(s_{1},f_{*}(s_{1}),U(f_{*}))$ for $s_{1}\in S.$ This fact and the proof of Theorem 1 or 2 imply that $v_{*}(s)=U(f_{*})(s)$ for all $s\in S.$ Let $\pi=(\pi_{n})_{n\in\mathbb{N}}$ be any Markov policy. We have $v_{*}(s)={\cal T}v_{*}(s)\geq{\cal T}_{\pi_{n}}v_{*}(s)$ for all $n\in\mathbb{N}$ and $s\in S.$ Iterating these inequalities we obtain for any $m\in\mathbb{N}$ that
$v_{*}(s)$$\geq$ ${\cal T}_{\pi_{1}}v_{*}(s_{1})\geq{\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}v_{*}(s_{1})\geq\cdots\geq{\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{m}}v_{*}(s_{1})$
$\geq$ ${\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{m}}\hat{0}(s_{1})=U_{m}(\pi)(s_{1}).$
Hence, it follows that
$U(f_{*})(s_{1})=v_{*}(s_{1})\geq\lim_{m\to\infty}U_{m}(\pi)(s_{1})=U(\pi)(s_{1})$
for every $s_{1}\in S.$Box
Case (B): $\gamma>1$ and $\rho>1.$ Moreover, (B1)-(B2) are satisfied and if $\theta\in(0,1),$ then (B3) holds and if $\theta>1,$ then (B4) holds.
For any $f\in\Phi$ we define ${\cal T}_{f}\hat{0}(s):=r(s,f(s))^{\frac{1}{1-\rho}}$ for $s\in S.$ For $\psi>0$ we put ${\cal T}_{f}\psi(s):=H(s,f(s),\psi).$ Then for any Markov policy $\pi=(\pi_{n})_{n\in\mathbb{N}}\in\Pi_{M}$, the recursive utility in the finite time horizon $U_{n}(\pi)(s_{1})$ can be defined as in (12), i.e.,
$U_{n}(\pi)(s_{1})={\cal T}_{\pi_{1}}{\cal T}_{\pi_{2}}\cdots{\cal T}_{\pi_{n}}\hat{0}(s_{1}),\ \ n\in\mathbb{N}.$
It should be noted that if $01-\gamma<$ 0 and $1-\rho<0,$ then we have $U_{m+1}(\pi)\leq U_{m}(\pi).$ Therefore, we can define
$U(\pi)(s_{1}):=\lim_{m\to\infty}U_{m}(\pi)(s_{1}).$
The sequence ${\cal T}^{m}\hat{0}$ of $m$ iterations of the operator ${\cal T}$ defined in (14) with itself is non-increasing. Under our assumptions ${\cal T}^{n}\hat{0}(\cdot)$ is continuous on $S$ for every $n\in\mathbb{N}.$ Let
$w_{0}(s):=\lim_{m\to\infty}{\cal T}^{m}\hat{0}(s)\quad\mbox{for }s\in S.$
The function $w_{0}$ is upper semicontinuous and using Proposition 10.1 in Schäl (1975), we infer that for all $s\in S$ it holds
$w_{0}(s)$ $=$ $\lim_{m\to\infty}{\cal T}^{m}\hat{0}(s)=\lim_{m\to\infty}\max_{a\in A(s)}H(s,a,{\cal T}^{m-1}\hat{0})=\max_{a\in A(s)}\lim_{m\to\infty}H(s,a,{\cal T}^{m-1}\hat{0})$ (15)
$=$ $\max_{a\in A(s)}H(s,a,w_{0}).$
Theorem 6
In Case (B), there exists a stationary optimal policy in the class of all policies.
Proof By Theorem 3 or 4 (depending on whether (B3) or (B4) is assumed), the Bellman equation has a unique solution $v_{*}$ and there exists $f_{*}\in\Phi$ such that
$v_{*}(s_{1})={\cal T}v_{*}(s_{1})=H(s_{1},f_{*}(s_{1}),v_{*})\quad\mbox{for all}\quad s_{1}\in S.$
Our arguments used in the proofs of Theorems 3 and 4 and (15) imply that $w_{0}(s)=v_{*}(s)=U(f_{*})(s)$ for all $s\in S.$ Fix any state $s_{1}\in S$, a number $\epsilon>0$ and a Markov policy $\pi=(\pi_{n})_{n\in\mathbb{N}}.$ There exists $N>0$ such that
$w_{0}(s_{1})=v_{*}(s_{1})\geq{\cal T}^{m}\hat{0}(s_{1})-\epsilon\geq U_{m}(\pi)(s_{1})-\epsilon,\quad\mbox{for all}\quad m>N.$
Hence
$U(f_{*})(s_{1})=v_{*}(s_{1})\geq\lim_{m\to\infty}(U_{m}(\pi)(s_{1})-\epsilon)=U(\pi)(s_{1})-\epsilon.$
Since $\epsilon>0$ can be arbitrarily small, it follows that $U(f_{*})(s_{1})\geq U(\pi)(s_{1}).$Box
3 Proofs of Theorems 1-4
The proof of Theorem 1 is proceeded by an auxiliary lemma.
Lemma 1
Under assumptions (A1)-(A3), the assertions $(a)$-$(b)$ of Theorem 1 hold.
Proof   First, we must show that $F_{1}$ maps ${\cal U}_{+}^{\omega}$ into itself. Let $V\in{\cal U}_{+}^{\omega}.$ Note that from (A2) and (A3), we get
00 0 $\leq$ $F_{1}V(s)\leq\left[M\omega(s)+\beta\left(\int_{S}\|V\|_{\omega}\omega(s^{\prime})q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]^{\theta}$
$\leq$ $\left[M\omega(s)^{\frac{1}{\theta}}+\beta\left(c\|V\|_{\omega}\omega(s)\right)^{\frac{1}{\theta}}\right]^{\theta}\leq\omega(s)\left[M+\beta c^{\frac{1}{\theta}}\|V\|_{\omega}^{\frac{1}{\theta}}\right]^{\theta},\quad s\in S.$
Additionally, by Lemma 8.5.5 in Hernández-Lerma and Lasserre (1999), it follows that the function
$(s,a)\to\int_{S}V(s^{\prime})q(ds^{\prime}|s,a)$
is upper semicontinuous. Therefore, $(s,a)\to H_{1}(s,a,V)$ is upper semicontinuous. Consequently, by (A1) and Lemma 3.8.3(c) in Hernández-Lerma and Lasserre (1999) the function $F_{1}V$ is upper semicontinuous on $S.$ Next we prove that $F_{1}$ is a $c\beta^{\theta}$-contraction. Let $V_{1},V_{2}\in{\cal U}^{\omega}_{+}$. Consider the norm
$\|\boldsymbol{z}\|_{\frac{1}{\theta}}=\left(|z_{1}|^{\frac{1}{\theta}}+|z_{2}|^{\frac{1}{\theta}}\right)^{\theta}$
for any vector $\boldsymbol{z}=(z_{1},z_{2})\in{\mathbb{R}}^{2}.$ Observe that $\|\cdot\|_{\frac{1}{\theta}}$ is indeed the norm, since $\frac{1}{\theta}>1.$ Now take
$\boldsymbol{x}:=\left(r^{\theta},\beta^{\theta}t_{1}\right),\quad\boldsymbol{y}:=\left(r^{\theta},\beta^{\theta}t_{2}\right),$
where we use the following abbreviations
$r^{\theta}:=r(s,a)^{\theta},\quad t_{i}:=\int_{S}V_{i}(s^{\prime})q(ds^{\prime}|s,a)$
for $(s,a)\in\mathbb{K}.$ We notice that
$\|\boldsymbol{x}\|_{\frac{1}{\theta}}=H_{1}(s,a,V_{1})\ \mbox{ and }\ \|\boldsymbol{y}\|_{\frac{1}{\theta}}=H_{1}(s,a,V_{2})$
with $H_{1}$ in (6). Although for convenience of notation we do not indicate dependence of $\boldsymbol{x}$ and $\boldsymbol{y}$ on $(s,a)$ we keep it in mind. Using these facts, we obtain
$|F_{1}V_{1}(s)-F_{1}V_{2}(s)|$$\leq$ $\max_{a\in A(s)}|H_{1}(s,a,V_{1})-H_{1}(s,a,V_{2})|=\max_{a\in A(s)}\left|\|\boldsymbol{x}\|_{\frac{1}{\theta}}-\|\boldsymbol{y}\|_{\frac{1}{\theta}}\right|$
$\leq$ $\max_{a\in A(s)}\|\boldsymbol{x}-\boldsymbol{y}\|_{\frac{1}{\theta}}=\max_{a\in A(s)}\|(0,\beta^{\theta}(t_{1}-t_{2}))\|_{\frac{1}{\theta}}$
$=$ $\max_{a\in A(s)}\left(0+\left|\beta^{\theta}(t_{1}-t_{2})\right|^{\frac{1}{\theta}}\right)^{\theta}=\max_{a\in A(s)}\beta^{\theta}|t_{1}-t_{2}|$
$\leq$ $\max_{a\in A(s)}\beta^{\theta}\int_{S}\frac{|V_{1}(s^{\prime})-V_{2}(s^{\prime})|}{\omega(s^{\prime})}\omega(s^{\prime})q(ds^{\prime}|s,a)$
$\leq$ $\beta^{\theta}c\|V_{1}-V_{2}\|_{\omega}\omega(s),\quad s\in S.$
Now dividing all sides of these inequalities by $\omega(s)$ and taking the supremum over all $s\in S$ we get
$\|F_{1}V_{1}-F_{1}V_{2}\|_{\omega}\leq c\beta^{\theta}\|V_{1}-V_{2}\|_{\omega}.$
Hence, $F_{1}$ is a $c\beta^{\theta}$-contraction. Since ${\cal U}_{+}^{\omega}$ is a closed subset of the Banach space ${\cal C}^{\omega}$ we conclude from the Banach contraction principle that there a unique $w_{*}\in{\cal U}^{\omega}_{+}$ such that
$w_{*}(s)=F_{1}w_{*}(s)\quad\mbox{for }s\in S.$
Putting $v_{*}:=w_{*}^{\frac{1}{1-\gamma}}$ we get that $v_{*}$ is non-negative, upper semicontinuous and
$v_{*}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}v_{*}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}}$ (16)
for all $s\in S.$ To finish this proof note that (7) now follows from (1) with $F=F_{1}.$Box
Proof of Theorem 1  In view of Lemma 1 it remains to prove assertions $(c)$ and $(d).$ The existence of a maximiser $f_{*}\in\Phi$ of the right-hand side of (16) follows from our assumptions (A1)-(A3) and Corollary 1 in Brown and Purves (1973). Let ${\cal B}^{\omega}$ be the set of all Borel measurable functions on $S$ for which $\|\cdot\|_{\omega}$-norm is finite and ${\cal B}_{+}^{\omega}$ be the set of non-negative functions in ${\cal B}^{\omega}.$ For any $f\in\Phi$ and $w\in{\cal B}_{+}^{\omega}$, we define the operator $F_{1f}$ as follows
$F_{1f}w(s):=H_{1}(s,f(s),w),\quad s\in S.$
Using the measurability results of Brown and Purves (1973), similarly as above, we can show that $F_{1f}w\in{\cal B}_{+}^{\omega}.$ Moreover, $F_{1f}$ is a $c\beta^{\theta}$-contraction. Therefore, the operator $F_{1f}$ has s a unique fixed point $w_{f}\in{\cal B}_{+}^{\omega}$ and for any function $w_{0}\in{\cal B}_{+}^{\omega}$
$\lim_{n\to\infty}\|F^{n}_{1f}w_{0}-w_{f}\|_{\omega}=0.$
Observe also that $F_{1f}$ is monotone, i.e. $F_{1f}w_{1}\geq F_{1f}w_{2}$ whenever $w_{1}\geq w_{2}$ and $w_{1},w_{2}\in{\cal B}_{+}^{\omega}.$ For any $f\in\Phi$, we have $F_{1f}w_{*}\leq w_{*}.$ By iteration of this inequality we get $F^{n}_{1f}w_{*}\leq w_{*}$ and consequently $w_{f}\leq w_{*},$ since $w_{*}\in{\cal U}_{+}^{\omega}\subset{\cal B}_{+}^{\omega}.$ On the other hand, due to Lemma 1 we have $F_{1f_{*}}w_{*}=w_{*}=F_{1}w_{*}.$ Again iterating $F_{1f_{*}}$ we obtain $\lim_{n\to\infty}F^{n}_{1f_{*}}w_{*}=w_{f_{*}}=w_{*}.$ Summing up,
$w_{f}\leq w_{*}=w_{f_{*}}\quad\mbox{for all }\ f\in\Phi,$
which implies via transformation $v:=w^{\frac{1}{1-\gamma}}$ with $01-\gamma>$ 0 that
$v_{f}\leq v_{*}=v_{f_{*}}\quad\mbox{for all }\ f\in\Phi.$
This means that $f_{*}$ is an optimal stationary policy. $\Box$
The proof of Theorem 2 is proceeded by an auxiliary fact.
Lemma 2
Under assumptions (A1)-(A2) and (A4), the assertions $(a)$-$(b)$ of Theorem 2 hold.
Proof   First, we must show that $F_{2}:{\cal U}_{+}^{\omega}\to{\cal U}_{+}^{\omega}.$ Indeed, note that from (A2) and (A4), we get
$0\leq|F_{2}V(s)|$$\leq$ $\left[M\omega(s)+\beta\left(\int_{S}\|V\|^{\theta}_{\omega}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]$
$\leq$ $\left[M\omega(s)+\beta\left(c\|V\|^{\theta}_{\omega}\omega(s)^{\theta}\right)^{\frac{1}{\theta}}\right]\leq\omega(s)\left[M+\beta c^{\frac{1}{\theta}}\|V\|_{\omega}\right].$
Again, by Lemma 8.5.5 in Hernández-Lerma and Lasserre (1999), it follows that the function
$(s,a)\to\int_{S}V(s^{\prime})q(ds^{\prime}|s,a)$
is upper semicontinuous. Therefore, $(s,a)\to H_{2}(s,a,V)$ is upper semicontinuous. Consequently, by (A1) and the Maximum Theorem of Berge (see Section 17.5 in Aliprantis and Border (2006)), the function $F_{2}V$ is upper semicontinuous on $S.$ Now we prove that $F_{2}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Let $V_{1},V_{2}\in{\cal U}^{\omega}_{+}$. Then, by the Minkowski inequality we have
$|F_{2}V_{1}(s)-F_{2}V_{2}(s)|$$\leq$ $\max_{a\in A(s)}\beta\left|\left(\int_{S}V_{1}(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}-\left(\int_{S}V_{2}(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right|$
$\leq$ $\max_{a\in A(s)}\beta\left(\int_{S}|V_{1}(s^{\prime})-V_{2}(s^{\prime})|^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}$
$\leq$ $\max_{a\in A(s)}\beta\left(\int_{S}\|V_{1}-V_{2}\|_{\omega}^{\theta}\omega(s^{\prime})^{\theta}q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}$
$\leq$ $\beta\|V_{1}-V_{2}\|_{\omega}c^{\frac{1}{\theta}}\omega(s),\quad s\in S.$
Now dividing all sides by $\omega(s)$ and taking the supremum over all $s\in S$ we get
$\|F_{2}V_{1}-F_{2}V_{2}\|_{\omega}\leq c^{\frac{1}{\theta}}\beta\|V_{1}-V_{2}\|_{\omega}.$
Hence, $F$ is $c^{\frac{1}{\theta}}\beta$-contraction. Since ${\cal U}_{+}^{\omega}$ is a closed subset of the Banach space ${\cal C}^{\omega}$ we conclude from the Banach contraction principle that there exists $w_{*}\in{\cal U}^{\omega}_{+}$ such that
$w_{*}(s)=F_{2}w_{*}(s)\quad s\in S.$
Putting $v_{*}:=w_{*}^{\frac{1}{1-\rho}}$ we get that
$v_{*}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}v_{*}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},\quad s\in S.$ (17)
for all $s\in S.$ Note that (8) follows from (1) with $F=F_{2}.$Box
Proof of Theorem 2  In view of Lemma 2 it remains to prove assertions $(c)$ and $(d).$ The existence of a maximiser $f_{*}\in\Phi$ of the right-hand side of (17) follows from our assumptions (A1)-(A2), (A4) and a measurable selection theorem, see Corollary 1 in Brown and Purves (1973). Again we consider the space ${\cal B}_{+}^{\omega}.$ For any $f\in\Phi$ and $w\in{\cal B}_{+}^{\omega}$ we define the operator $F_{2f}$ as follows
$F_{2f}w(s):=H_{2}(s,f(s),w),\quad s\in S.$
Similarly as in the proof of Theorem 1 we may show that $F_{2f}w\in{\cal B}_{+}^{\omega}.$ Moreover, $F_{2f}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Therefore, the operator $F_{2f}$ has s a unique fixed point $w_{f}\in{\cal B}_{+}^{\omega}$ and for any function $w_{0}\in{\cal B}_{+}^{\omega}$
$\lim_{n\to\infty}\|F^{n}_{2f}w_{0}-w_{f}\|_{\omega}=0.$
Observe also that $F_{2f}$ is monotone, i.e. $F_{2f}w_{1}\geq F_{2f}w_{2}$ whenever $w_{1}\geq w_{2}$ and $w_{1},w_{2}\in{\cal B}_{+}^{\omega}.$ For any $f\in F$ we have $F_{2f}w_{*}\leq w_{*}.$ By iteration of this inequality we get $F^{n}_{2f}w_{*}\leq w_{*}$ and consequently $w_{f}\leq w_{*}.$ On the other hand, $F_{2f_{*}}w_{*}=w_{*}=F_{2}w_{*}.$ Again iterating $F_{2f_{*}}$ we obtain $\lim_{n\to\infty}F^{n}_{2f_{*}}w_{*}=w_{f_{*}}=w_{*}.$ Thus,
$w_{f}\leq w_{*}=w_{f_{*}}\quad\mbox{for all }\ f\in\Phi,$
which via transformation $v:=w^{\frac{1}{1-\rho}}$ with $01-\rho>$ 0 gives that
$v_{f}\leq v_{*}=v_{f_{*}}\quad\mbox{for all }\ f\in\Phi.$
This means that $f_{*}$ is an optimal stationary policy. $\Box$
Proof of Theorem 3  Proceeding as in Lemma 2 we obtain that $F_{3}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Moreover, from (B2) and (B4) for any $V\in{\cal C}_{+}^{\omega}$ we have
$0<\|F_{3}V\|_{\omega}\leq M+\beta c^{\frac{1}{\theta}}\|V\|_{\omega}.$
Note that Lemma 8.5.5 in Hernández-Lerma and Lasserre (1999), (B1)-(B2), (B4) and the Maximum Theorem of Berge (Section 17.5 in Aliprantis and Border (2006)) imply that $F_{3}V$ is continuous on $S.$ Since ${\cal C}_{+}^{\omega}$ is a closed subset of the Banach space ${\cal C}^{\omega}$ we conclude from the Banach contraction principle that there exists $w_{*}\in{\cal C}^{\omega}_{+}$ such that
$w_{*}(s)=F_{3}w_{*}(s)\quad s\in S.$
Note that $w_{*}(s)\geq\min_{a\in A(s)}r(s,a)>0$ for every $s\in S.$ Putting $v_{*}:=w_{*}^{\frac{1}{1-\rho}}$ we get that
$v_{*}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}v_{*}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},\quad s\in S.$ (18)
for all $s\in S.$ Clearly, (9) follows from (1) with $F=F_{3}.$ Now, we prove (c) and (d). Let $f_{*}\in\Phi$ be a maximiser of the right-hand side of (18). For any $f\in\Phi$ and $w\in{\cal B}_{+}^{\omega}$ set
$F_{3f}w(s):=H_{3}(s,f(s),w),\quad s\in S.$
It follows that $F_{3f}w\in{\cal B}_{+}^{\omega}$ and $F_{3f}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Therefore, the operator $F_{3f}$ has s a unique fixed point $w_{f}\in{\cal B}_{+}^{\omega}$ and for any function $w_{0}\in{\cal B}_{+}^{\omega}$
$\lim_{n\to\infty}\|F^{n}_{3f}w_{0}-w_{f}\|_{\omega}=0.$
Additionally, $F_{3f}$ is also monotone, i.e. $F_{3f}w_{1}\geq F_{3f}w_{2}$ if $w_{1}\geq w_{2}$ and $w_{1},w_{2}\in{\cal B}_{+}^{\omega}.$ For any $f\in F$ we have $F_{3f}w_{*}\geq w_{*}.$ By iteration of this inequality we get $F^{n}_{3f}w_{*}\geq w_{*}$ and consequently $w_{f}\geq w_{*}.$ On the other hand, $F_{3f_{*}}w_{*}=w_{*}=F_{3}w_{*}.$ Again iterating $F_{3f_{*}}$ we obtain $\lim_{n\to\infty}F^{n}_{3f_{*}}w_{*}=w_{f_{*}}=w_{*}.$ Thus,
$w_{f}\geq w_{*}=w_{f_{*}}\quad\mbox{for all }\ f\in\Phi,$
which via transformation $v:=w^{\frac{1}{1-\rho}}$ with $01-\rho<$ 0 yields that
$v_{f}\leq v_{*}=v_{f_{*}}\quad\mbox{for all }\ f\in\Phi.$
This means that $f_{*}$ is an optimal stationary policy. $\Box$
Proof of Theorem 4  Similarly as in Lemma 1 we show that $F_{4}$ is $c\beta^{\theta}$-contraction. Moreover, from (B2) and (B3) for any $V\in{\cal C}_{+}^{\omega}$ we have
$0<\|F_{4}V\|_{\omega}\leq\left[M+\beta c^{\frac{1}{\theta}}\|V\|_{\omega}^{\frac{1}{\theta}}\right]^{\theta}.$
Lemma 8.5.5 in Hernández-Lerma and Lasserre (1999), (B1)-(B3) and the Maximum Theorem of Berge (Section 17.5 in Aliprantis and Border (2006)) imply the continuity of $F_{4}V$ on $S.$ From the Banach contraction principle there exists $w_{*}\in{\cal C}^{\omega}_{+}$ such that
$w_{*}(s)=F_{4}w_{*}(s)\quad s\in S.$
By (B2) $w_{*}(s)\geq\min_{a\in A(s)}r(s,a)>0$ for every $s\in S.$ Putting $v_{*}:=w_{*}^{\frac{1}{1-\gamma}}$ we get that
$v_{*}(s)=\max_{a\in A(s)}\left[r(s,a)+\beta\left(\int_{S}v_{*}(s^{\prime})^{1-\gamma}q(ds^{\prime}|s,a)\right)^{\frac{1-\rho}{1-\gamma}}\right]^{\frac{1}{1-\rho}},\quad s\in S.$ (19)
for all $s\in S.$ Obviously, (10) follows from (1) with $F=F_{4}.$ It remains to prove (c) and (d). Let $f_{*}\in\Phi$ be a maximiser of the right-hand side of (19). For $f\in\Phi$ and $w\in{\cal B}_{+}^{\omega}$ set
$F_{4f}w(s):=H_{4}(s,f(s),w),\quad s\in S.$
Then, $F_{4f}w\in{\cal B}_{+}^{\omega}$ and $F_{4f}$ is $c^{\frac{1}{\theta}}\beta$-contraction. Hence, the operator $F_{4f}$ has s a unique fixed point $w_{f}\in{\cal B}_{+}^{\omega}$ and for any function $w_{0}\in{\cal B}_{+}^{\omega}$
$\lim_{n\to\infty}\|F^{n}_{4f}w_{0}-w_{f}\|_{\omega}=0.$
Additionally, $F_{4f}$ is also monotone, i.e. $F_{4f}w_{1}\geq F_{4f}w_{2}$ if $w_{1}\geq w_{2}$ and $w_{1},w_{2}\in{\cal B}_{+}^{\omega}.$ For any $f\in F$ we have $F_{4f}w_{*}\geq w_{*}.$ By iteration we obtain $F^{n}_{4f}w_{*}\geq w_{*}$ and finally $w_{f}\geq w_{*}.$ On the other hand, $F_{4f_{*}}w_{*}=w_{*}=F_{4}w_{*}.$ Iterating $F_{4f_{*}}$ we get $\lim_{n\to\infty}F^{n}_{4f_{*}}w_{*}=w_{f_{*}}=w_{*}.$ Thus,
$w_{f}\geq w_{*}=w_{f_{*}}\quad\mbox{for all }\ f\in\Phi,$
which via transformation $v:=w^{\frac{1}{1-\gamma}}$ with $01-\gamma<$ 0 yields that
$v_{f}\leq v_{*}=v_{f_{*}}\quad\mbox{for all }\ f\in\Phi.$
This proves that $f_{*}$ is an optimal stationary policy. $\Box$
4 Remarks on an application of Du’s fixed point theorem and Banach’s contraction principle
In this section, we consider bounded utility functions.
4.1 Du’s theorem and convex/concave operators
[…]
Proposition 1
Let $T:I\to I$ be a monotone increasing operator. If either (i) $T$ is convex on $I$ and for some $\epsilon\in(0,1)$
$Tg_{2}\leq(1-\epsilon)g_{2}+\epsilon g_{1},$ (20)
or (ii) $T$ is concave on $I$ and for some $\epsilon\in(0,1)$
$Tg_{1}\geq(1-\epsilon)g_{1}+\epsilon g_{2},$ (21)
then $T$ has a unique fixed point $v_{*}\in I.$ Moreover, for any $v_{0}\in I$ and the sequence defined by $v_{n}=Tv_{n-1},$ $n\in\mathbb{N},$ there exists $B>0$ such that
$\|v_{n}-v_{*}\|\leq B(1-\epsilon)^{n},\ \ n\in\mathbb{N}.$ (22)
Since ${\cal C}_{+}$ is a normal cone in $\cal C$ with constant $N=1$, from the proofs of the theorem of Du (1990) or Theorem 2.1.2 in Zhang (2012), it follows that, if $T$ satisfies $(i),$ then
$B=\|g_{1}-g_{2}\|+\frac{2\|Tg_{2}-g_{2}\|}{\epsilon^{2}}$ (23)
and if $T$ satisfies $(ii),$ then
$B=\|g_{1}-g_{2}\|+\frac{2\|Tg_{1}-g_{1}\|}{\epsilon^{2}}.$ (24)
Remark 6
Inequalities (20) and (21) are called boundary conditions. They are crucial in proving the geometric convergence of iterations of the operator $T$ to a fixed point that belongs to the interior of the set $I.$
Below we consider two cases discussed in Ren and Stachurski (2020) who applied the transformation $v=w^{1-\gamma}$. Under this transformation $H$ defined in (2) can be written as
${\cal H}(s,a,v)=\left[r(s,a)+\beta\left(\int_{S}v(s^{\prime})q(ds^{\prime}|s,a)\right)^{\frac{1}{\theta}}\right]^{\theta},\quad(s,a)\in\mathbb{K}.$
Recall also that $\theta=\frac{1-\gamma}{1-\rho}.$
Case 1: $0<\rho<\gamma<1,$ i.e., $\theta\in(0,1).$ In this setting, Bellman equation (5) is equivalent to
$v(s)=Tv(s):=\max_{a\in A(s)}{\cal H}(s,a,v),$ (25)
From the convexity of the function $\psi(t)=(b+\beta t^{1/\theta})^{\theta}$ on $\mathbb{R}_{+}$ for any $b\geq 0$, it follows that for each $v_{1}$, $v_{2}$ and $\xi\in(0,1)$,
${\cal H}(s,a,\xi v_{1}+(1-\xi)v_{2})\leq\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2}).$ (26)
Hence,
$T(\xi v_{1}(s)+(1-\xi)v_{2})(s)\leq\xi Tv_{1}(s)+(1-\xi)Tv_{2}(s),$
i.e., $T$ is convex.
Case 3: $1<\rho<\gamma,$ i.e., $\theta>1.$ In this setting, Bellman equation (5) is equivalent to
$v(s)=Tv(s):=\min_{a\in A(s)}{\cal H}(s,a,v),\quad s\in S.$ (27)
Since the function $\psi(t)=(b+\beta t^{1/\theta})^{\theta}$ is concave on $\mathbb{R}_{+}$ for any $b\geq 0,$ it follows that for each $v_{1}$, $v_{2}$ and $\xi\in(0,1)$,
${\cal H}(s,a,\xi v_{1}+(1-\xi)v_{2})\geq\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2}).$ (28)
Hence,
$T(\xi v_{1}(s)+(1-\xi)v_{2})(s)\geq\xi Tv_{1}(s)+(1-\xi)Tv_{2}(s),$
i.e., $T$ is concave.
Below we comment on two other cases for which the Banach contraction mapping theorem is applied in Subsection 2.3 after suitable transformations. We explain why the fixed point theorem of Du (1990) cannot be applied.
Case 2: $0<\gamma<\rho<1,$ i.e., $\theta>1.$ Now Bellman equation (5) is equivalent to (25). Now $v\to{\cal H}(s,a,v)$ is concave. Therefore, (28) holds for each $v_{1}$, $v_{2}$ and $\xi\in(0,1).$ Hence,
$\max_{a\in A(s)}{\cal H}(s,a,\xi v_{1}+(1-\xi)v_{2})\geq\max_{a\in A(s)}(\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2})),$
but
$\max_{a\in A(s)}(\xi{\cal H}(s,a,v_{1})+(1-\xi{\cal H}(s,a,v_{2}))\not\geq\xi\max_{a\in A(s)}{\cal H}(s,a,v_{1})+(1-\xi)\max_{a\in A(s)}{\cal H}(s,a,v_{2}).$
Thus, in this case $T$ is not concave and the theorem of Du (1990) cannot be applied.
Case 4: $1<\gamma<\rho,$ i.e., $\theta\in(0,1).$ Here, Bellman equation (5) is equivalent to (27). The function $\psi(t)=(b+\beta t^{1/\theta})^{\theta}$ is convex on $\mathbb{R}_{+}$ and therefore, (26) holds for each $v_{1}$, $v_{2}$ and $\xi\in(0,1).$ Hence,
$T(\xi v_{1}(s)+(1-\xi)v_{2})(s)\leq\min_{a\in A(s)}\left(\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2})\right)$
but
$\min_{a\in A(s)}\left(\xi{\cal H}(s,a,v_{1})+(1-\xi){\cal H}(s,a,v_{2})\right)\not\leq\xi\min_{a\in A(s)}{\cal H}(s,a,v_{1})+(1-\xi)\min_{a\in A(s)}{\cal H}(s,a,v_{2}).$
Thus, $T$ is not convex and the theorem of Du (1990) cannot be applied to this case.
4.2 Comparing the speed of convergence in Du’s theorem and Banach’s theorem
Our aim in this section is to compare the iteration algorithms proposed by Ren and Stachurski (2020, 2021) with our modifications. We are interested in estimating the speed of convergence in both cases.
We remind that from the proof of the Banach contraction mapping principle, it follows that, if $F:{\cal C}_{+}\to{\cal C}_{+}$ is a $\delta$-contraction, then $F$ has a unique fixed point $w_{*}\in{\cal C}_{+}$ and for each $n\in\mathbb{N}$
$\|w_{n}-w_{*}\|\leq L\delta^{n}\quad\mbox{where}\quad L=\frac{\|F\hat{0}-\hat{0}\|}{1-\delta},$ (29)
$w_{0}\equiv\hat{0}$ and $w_{n}=Fw_{n-1},$ $n\in\mathbb{N}.$
Case 1: $10<\rho<\gamma<$ 1 . We assume that
$m=\min_{(s,a)\in\mathbb{K}}r(s,a)=0\quad\mbox{and}\quad M=\max_{(s,a)\in\mathbb{K}}r(s,a).$
We fix $y>0$ and similarly as in Ren and Stachurski (2020) we consider the constant functions
$g_{1}(s):=\hat{0}\quad\mbox{and}\quad g_{2}(s):=\left(\frac{M+y}{1-\beta}\right)^{\theta},\ \ s\in S.$
As noted by Ren and Stachurski (2020), the operator $T$ considered in this case is convex. Thus, the geometric convergence of iterations of $T$ with $v_{0}=g_{1}=\hat{0}$ can be analysed using the formula (23). Below we try to estimate the values of $B$ and $1-\epsilon$ when $\theta=1/2.$
Example 1
Assume that $M=1$, $\beta=0.9,$ $\rho=1/2$ and $\gamma=3/4.$ Then $\theta=1/2.$ Let
$\min_{s\in S}\max_{a\in A(s)}r(s,a)=0.$
We have
$\|g_{1}-g_{2}\|=\sqrt{10(1+y)}\quad\mbox{and}\quad Tg_{2}(s)=\sqrt{\max_{a\in A(s)}r(s,a)+9(1+y)}$
and consequently,
$\|Tg_{2}-g_{2}\|$ $=$ $\sqrt{10+10y}-\sqrt{9(1+y)+\min_{s\in S}\max_{a\in A(s)}r(s,a)}$
$=$ $\sqrt{10+10y}-3\sqrt{1+y}=(\sqrt{10}-3)\sqrt{1+y}.$
It is easy to see that (20) holds if
$\sqrt{1+9(1+y)}\leq(1-\epsilon)\sqrt{10+10y}.$
Hence,
$\epsilon\leq\epsilon(y):=1-\sqrt{\frac{1+0.9y}{1+y}},\quad y>0.$
Then
$1-\epsilon(y)=\sqrt{\frac{1+0.9y}{1+y}}.$
To obtain the fastest geometric convergence of iterations of the operator $T$ with starting function $v_{0}=\hat{0}$, we apply the following method. We can minimise $B(y)(1-\epsilon(y))$ where $B(y)=B$ is given in (23). Note that
$B(y)=B=\sqrt{10+10y}+\frac{2(\sqrt{10}-3)\sqrt{1+y}}{\left(1-\sqrt{\frac{1+0.9y}{1+y}}\right)^{2}}$
and
$B(y)(1-\epsilon(y))=\left(\sqrt{10}+\frac{2(\sqrt{10}-3)}{\left(1-\sqrt{\frac{1+0.9y}{1+y}}\right)^{2}}\right)\sqrt{1+0.9y}$
where $y>0.$ The minimum value of $B(y)(1-\epsilon(y))$ over the interval $(0,\infty)$ is attained at $y_{0}\approx 4.10707$ and equals $B(y_{0})(1-\epsilon(y_{0}))\approx 424.197.$ Next $1-\epsilon(y_{0})\approx 0.958947704>\sqrt{\beta}=\sqrt{0.9}\approx 0.948683298$ The reason for these results is that $T$ is not treated as a contraction mapping in Proposition 1. We note that
$\inf_{y>0}(1-\epsilon(y))=\lim_{y\to\infty}(1-\epsilon(y))=\sqrt{\beta}=\sqrt{0.9},$
but $\lim_{y\to\infty}B(y)=\infty.$ Thus, minimising only the expression $1-\epsilon(y)$ makes no sense.
The operator $F_{1}$ proposed in Subsection 2.3 to study iterations with starting function $v_{0}=\hat{0}$ is as $T$ in (25) with $\theta=1/2.$ We have shown that $F_{1}$ is a $\sqrt{\beta}$-contraction. Since the algorithm using the operator $F_{1}$ starts with $v_{0}=\hat{0},$ the constant $L$ in (29) is equal to 1/($19.48681/(1-\sqrt{\beta})=1/(1-\sqrt{0.9})=$ 19.4868 . It is obvious that the speed of convergence of iterations of the operator $F_{1}$ viewed as a $\sqrt{\beta}$-contraction is much better than in the former case. It should be noted that the solution to Bellman equation (5) in this example is $w_{*}^{4}.$
Case 3: $1<\rho<\gamma.$ We now assume that
$0<m=\min_{(s,a)\in\mathbb{K}}r(s,a)\quad\mbox{and}\quad M=\max_{(s,a)\in\mathbb{K}}r(s,a).$
We fix $x\in(0,m)$ and as in Ren and Stachurski (2020) we consider the constant functions
$g_{1}(s):=\left(\frac{x}{1-\beta}\right)^{\theta}\quad\mbox{and}\quad g_{2}(s):=\left(\frac{M}{1-\beta}\right)^{\theta}.$
As noted by Ren and Stachurski (2020), the operator $T$ considered in this case is concave. Therefore, the geometric convergence of iterations of $T$ with $v_{0}=g_{1}$ can be analysed using the formula (22) with $B$ defined in (24). Below we try to estimate $B$ and $1-\epsilon$ in (24) when $\theta=2.$
Example 2
Assume that $m=1,$ $M=5$ and
$\max_{s\in S}\min_{a\in A(s)}r(s,a)=3.$
Let $\beta=0.9,$ $\rho=1.25$ and $\gamma=1.5.$ Then $\theta=2$ and $x\in(0,1).$ Moreover,
$\|g_{1}-g_{2}\|=100(25-x^{2})\quad\mbox{and}\quad Tg_{1}(s)=\min_{a\in A(s)}(r(s,a)+9x)^{2}\leq(3+9x)^{2}$
for all $s\in S.$ Hence,
$\|Tg_{1}-g_{1}\|=(3+9x)^{2}-100x^{2}=9+54x-19x^{2}.$
One can easily see that (21) holds, if
$(1+9x)^{2}\geq(1-\epsilon)100x^{2}+\epsilon 2500.$
Thus, we have
$\epsilon\leq\epsilon(x):=\frac{1+18x-19x^{2}}{100(25-x^{2})},\quad x\in(0,1),$
and
$1-\epsilon(x)=\frac{2499-18x-81x^{2}}{100(25-x^{2})}.$
To get the best geometric convergence of iterations of the operator $T$ with starting function $v_{0}=g_{1}$, we can apply two approaches. First, we can try to minimise $B(x)(1-\epsilon(x)),$ where $B(x)=B$ is given in (24). We obtain
$B(x)=B=100(25-x^{2})+\frac{2\cdot 100^{2}(25-x^{2})^{2}(9+54x-19x^{2})}{(1+18x-19x^{2})^{2}},$
and
$B(x)(1-\epsilon(x))=(2499-18x-81x^{2})R(x)$
where
$R(x):=\left(1+\frac{200(25-x^{2})(9+54x-19x^{2})}{(1+18x-19x^{2})^{2}}\right),$
for $x\in(0,1).$ The minimum value of $B(x)(1-\epsilon(x))$ over the interval $(0,1)$ is attained at $x_{0}\approx 0.378544$ and is equal to $B(x_{0})(1-\epsilon(x_{0}))\approx 1.27143\times 10^{7}.$ This is a surprisingly large number. Moreover, $0.91-\epsilon(x_{0})\approx 0.997951789>\beta=$ 0.9 . The reason for such results is that $T$ is not a contraction mapping. Alternatively, one can find $x_{1}\in(0,1)$ that minimises $1-\epsilon(x)$ over $(0,1)$ and then calculate $B(x_{1}).$ Then
$x_{1}=\frac{79-8\sqrt{94}}{3}\approx 0.47904076,\quad\mbox{and}$
$B(x_{1})=\frac{3(377+4\sqrt{94})}{1250}\approx 0.997875453.$
Note that $1-\epsilon(x_{1})<1-\epsilon(x_{0}),$ but $B(x_{1})(1-\epsilon(x_{1}))>B(x_{0})(1-\epsilon(x_{0})).$ The operator $F_{3}$ proposed in Subsection 2.3 to study value iterations is
$F_{3}v(s)=\min_{a\in A(s)}\left(r(s,a)+\beta\sqrt{\int_{S}v(s^{\prime})^{2}q(ds^{\prime}|s,a)}\right)$
and is a $\beta$-contraction. Since the algorithm using the operator $F_{3}$ can start with $w_{0}=\hat{0},$ the constant $L$ in (29) is equal to $505/(1-\beta)=$ 50 . The difference between the algorithm using Du’s theorem for concave and increasing operators and our proposal is significant. One should remember that the solution to Bellman equation (5) in this example is the limit $\lim_{n\to\infty}w_{n}^{1/(1-\rho)}=w_{*}^{-4}.$
Remark 7
We would like to point out that in our method of studying existence and uniqueness of a solution to the Bellman equation we do not need boundary conditions mentioned in Remark 6. Therefore, we can drop the assumption made in the literature in case when $\gamma,\ \rho>1$ that the per-period utilities $u(s,a)\geq u_{0}$ for some $u_{0}>0$ and for all $(s,a)\in\mathbb{K}.$ However, if $\inf_{s\in S}\min_{a\in A(s)}u(s,a)=0,$ then $r(s,a)=(1-\beta)u(s,a)^{1-\rho}$ is unbounded and there should be a weight function $\omega$ satisfying extra assumptions.
5 Conclusions
We consider dynamic programming problems arising from Markov decision processes with the Epstein-Zin preferences involving certainty equivalents induced by power functions. The values of parameters $\gamma$ and $\rho$ are such that the aggregators that we deal with are according to Marinacci and Montrucchio (2010) of Thompson type. We have shown that in the two cases: (1) $10<\rho<\gamma<$ 1 and (2) $1<\rho<\gamma$ studied in Ren and Stachurski (2020, 2021) one can use the Banach contraction mapping principle instead of a fixed point theorem of Du (1990) for increasing and convex or concave operators acting on an ordered Banach space. The contraction mapping theorem is applied to some sets of non-negative functions defined on the state space, endowed with the standard weighted supremum metric. This is a much simpler way for studying the existence and uniqueness of recursive utilities for the models with the Epstein-Zin preferences than the method using the Thompson metrics as in Marinacci and Montrucchio (2010) or the fixed point theorem of Du (1990) as in Ren and Stachurski (2020, 2021). Our assumptions on the per-period utility functions are weaker than the corresponding ones in Marinacci and Montrucchio (2010) and Ren and Stachurski (2020, 2021). For example, we do not assume the boundary conditions assuring that the fixed point of the operator involved in the analysis lies in the interior of its domain. In dynamic programming the boundary conditions are connected with inequalities (20) and (21) in Proposition 1. The constant $\epsilon>0$ in (20) and (21) has an influence on the speed of convergence of the dynamic programming operator to a fixed point. Using the Banach theorem, we do not need to apply any boundary conditions and therefore, we are able to provide better estimation for the convergence of iterations of operators used in modified value iteration process. The relevance of this result is that our method requires a smaller number of iterations needed to reach a given accuracy of the solution to the Bellman equation than in the case of Du’s theorem. It seems that the third case considered by Ren and Stachurski (2020) where $0<\rho<1<\gamma$ cannot be analysed by the Banach contraction mapping principle. However, we wish to point out that our methodology applies to other two cases (1) $0<\gamma<\rho$ and (2) $1<\gamma<\rho,$ where the theorem of Du (1990) is inapplicable, since the dynamic programming operators are then neither convex nor concave.